### v3

# Shuffle-R1: Efficient RL framework for Multimodal Large Language Models via Data-centric Dynamic Shuffle

Abstract
Reinforcement learning (RL) has emerged as an effective post-training paradigm for enhancing the reasoning capabilities of multimodal large language model (MLLM). However, current RL pipelines often suffer from training inefficiencies caused by two underexplored issues: Advantage Collapsing, where most advantages in a batch concentrate near zero, and Rollout Silencing, where the proportion of rollouts contributing non-zero gradients diminishes over time. These issues lead to suboptimal gradient updates and hinder long-term learning efficiency. To address these issues, we propose Shuffle-R1, a simple yet principled framework that improves RL fine-tuning efficiency by dynamically restructuring trajectory sampling and batch composition. It introduces (1) Pairwise Trajectory Sampling, which selects high-contrast trajectories with large advantages to improve gradient signal quality, and (2) Advantage-based Batch Shuffle, which increases exposure of valuable rollouts through strategic batch reshuffling. Experiments across multiple reasoning benchmarks demonstrate that our framework consistently outperforms strong RL baselines with minimal computational overhead. These results highlight the importance of data-centric adaptations for more efficient RL training in MLLM.
footnotetext: 🖂 Corresponding author.
   This work was done during the research internship of Linghao Zhu, Yiran Guan, and Dingkang Liang at Xiaomi.
1 Introduction
Inspired by the slow and deliberate thinking process in human cognition, reasoning capabilities empower Large Language Model (LLM) to plan, reflect, and generalize beyond mere memorization (Chu et al. 2025a; Kang et al. 2024). Recent advances have demonstrated that integrating reinforcement learning (RL) into the training of LLMs can significantly enhance such capabilities, particularly in complex domains such as mathematical problem solving and code generation (Guo et al. 2025; Deepmind 2025; OpenAI 2024; Seed et al. 2025). Notably, DeepSeek-R1 (Guo et al. 2025) leverage reward signals derived exclusively from verifiable outcomes to yield impressive performance gains. Beyond textual tasks, RL has also seen increasing application in multimodal domains, including object detection, segmentation, and video understanding (Li et al. 2025; Liu et al. 2025c, b; Wang et al. 2025c), highlighting its potential to support generalizable reasoning across modalities.
Figure 1: (a): Advantage Collapsing, where most advantages concentrate near zero. (b): Rollout Silencing, where the ratio of rollouts with non-zero gradient consistently drops. The phenomenon gets worse in larger models.
Despite the growing interest in RL for LLM and Multimodal LLM (MLLM), several practical challenges remain unresolved, including slow convergence, training instability, and suboptimal efficiency. Recent studies have proposed various improvements, such as empirical optimization of training configurations (Yu et al. 2025), integration of experience replay (Wang et al. 2025a), and refinement of objective functions (Chu et al. 2025b). Other works focus on reward optimization (Liu et al. 2025a; Ma et al. 2025).
However, most of these approaches remain confined to the standard RL training paradigm, where trajectories are uniformly sampled without considering their varying informativeness or difficulty. This static strategy overlooks a crucial insight that not all learning signals are equally valuable, which leads us to consider a fundamental question: Can dynamically prioritizing trajectories provide richer gradient information and lead to more efficient training?
Motivated by this question, we examine the current RL training practices and identify two critical yet might underexplored inefficiencies. First, Advantage Collapsing emerges when most computed advantages cluster excessively near zero, drowning out informative signals from trajectories with large-magnitude advantages, resulting in weak or negligible gradient updates (Fig. 1(a)). Second, Rollout Silencing arises as the fraction of rollouts contributing non-zero gradients steadily declines during training (Fig. 1(b)), leading to wasted computation and underutilization of data. These two findings demonstrate a pressing need for adaptive mechanisms to prioritize, reuse, and reallocate gradient exposure toward informative samples.
In this paper, we introduce Shuffle-R1, an innovative framework designed to dynamically prioritize and amplify critical gradient information. Rather than asking how to design better rewards, we try to ask what data should the model actually update on during RL fine-tuning. Specifically, Shuffle-R1 introduces two elegantly simple yet highly effective modules: (1) Pairwise Trajectory Sampling, which selects high-contrast trajectory pairs with large-magnitude advantages from an extended rollout pool, concentrating learning signals to mitigate Advantage Collapsing; and (2) Advantage-based Batch Shuffle, which dynamically reshapes training batches to prioritize informative trajectories while down-weighting ineffective ones, alleviating Rollout Silencing to improve data utilization. Our framework is lightweight, modular, and seamlessly integrable with existing RL algorithms. Experiments demonstrate that Shuffle-R1 significantly improves model performance across challenging multimodal reasoning tasks, even surpassing GPT-4o and Claude-3.7 on MathVerse and MathVista. Moreover, it achieves competitive performance to GRPO while requiring only half of the training steps.
In summary, our contributions are three-fold: 1) We reveal two critical yet underexplored limitations that undermine training efficiency in RL fine-tuning for MLLM, i.e., Advantage Collapsing and Rollout Silencing. 2) We propose Shuffle-R1, a novel and adaptive RL framework that dynamically selects high-contrast trajectories and reshapes training batches to emphasize informative samples. 3) Extensive experiments across model scales and both in-domain and out-of-domain benchmarks demonstrate the effectiveness and generalizability of our framework.
2 Related Work
2.1 Large Reasoning Models
Researchers have explored various approaches to equip LLM with reasoning ability. Some early studies performed SFT on complex long chain-of-thought data, leading to performance gains on reasoning tasks (Muennighoff et al. 2025; Ye et al. 2025; Yue et al. 2024; Guo et al. 2024). However, some claim that SFT merely enables the model to memorize the format of reasoning steps and long chains of thought, without fully grasping the ability to reason independently (Chu et al. 2025a; Kang et al. 2024; Ye et al. 2024; Allen-Zhu and Li 2024). Some researchers control the model to generate structured chain-of-thought instead of free generation, achieving systematic step-by-step reasoning output (Xu et al. 2025; Wu et al. 2025; Thawakar et al. 2025) Other works attempted to use test-time scaling like Monte Carlo Tree Search (MCTS) (Yao et al. 2023; Zhang et al. 2024a; Du et al. 2024; Yao et al. 2024) to facilitate complex reasoning by actively extending the output of the model.
Recently, models such as OpenAI o1/o3 (OpenAI 2024), DeepSeek-R1 (Guo et al. 2025), Seed-Thinking (Seed et al. 2025), and Kimi-k1.5 (Team et al. 2025) utilized RL to enable the model to explore independently, stimulating reasoning ability. In particular, DeepSeek-R1-Zero directly conducted RL on pre-trained model without instruction fine-tuning, with verifiable outcome reward functions to replace reward models, achieving surprising reasoning ability. The training algorithms for RL are also constantly being optimized (Yu et al. 2025; Chu et al. 2025b). Our work focuses on a deeper investigation of the efficiency of RL training and puts forward an effective solution to improve both the efficiency and performance of RL training.
2.2 Reinforcement Learning for MLLM
Following the success of DeepSeek-R1, a series of studies have transplanted RL into the training of MLLM and downstream visual tasks, such as Open Vocabulary Object Detection (Liu et al. 2025c), Reasoning Segmentation (Liu et al. 2025b), Video Understanding (Li et al. 2025), Video Localization (Wang et al. 2025c), etc. These works mainly focus on the applicability of RL in downstream tasks. Some other works focusing on improving the general reasoning ability of MLLM have achieved performance improvement on reasoning tasks by collecting a large amount of high-quality data (Meng et al. 2025; Huang et al. 2025; Yang et al. 2025; Zhang et al. 2025; Peng et al. 2025; Tan et al. 2025). These works mainly focus on the organization of high-quality reasoning data and the balance between SFT and RL in the training process.
Some researchers who conduct in-depth research on the RL mechanism have optimized the RL training process from various aspects, including adding contrastive reward mechanism (Li et al. 2025), actively introducing reflection tokens during rollouts (Wang et al. 2025a), optimizing the RL objective function and gradient update mechanism (Chu et al. 2025b), introducing more diverse rollouts (Liu et al. 2025a; Yao et al. 2025), etc. The core objective of these works is to optimize the RL training process. In our work, we propose a novel training framework that introduces dynamic and adaptive selection and resampling to queries and rollouts, reshaping data distribution for better training efficiency and model performance.
3 Preliminaries
Reinforcement Learning in LLM.
Reinforcement learning is originally introduced as a post-training technique to align LLM outputs with human preference. Recently, it has been used in fine-tuning stage to improve the reasoning ability of LLMs. Policy gradient algorithm is the most commonly used RL method, whose objective function is to maximize the expectation of return the model gets from the environment. Given a query $q$, multiple independent sampled responses $O=\{o_{1},o_{2},...,o_{N}\}$ are sampled from the old policy model $o\sim\pi_{\theta^{\prime}}(q)$, with rollout size $N$. Subsequently, the corresponding rewards $R=\{r_{1},r_{2},...,r_{N}\}$ are obtained for each response, in our case they are calculated by verifiable reward functions. Advantages $\hat{A}=\{A^{\pi_{\theta}}(r_{1}),A^{\pi_{\theta}}(r_{2}),...,A^{\pi_{\theta}}(r_{N})\}$ are calculated to assess the improvement over the current policy model. Importance sampling and clipping are introduced to stabilize the training. The core objective function is defined as:
$\begin{split}&\mathcal{J}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta^{\prime}}(\cdot|q)}\\ &\frac{1}{\sum_{i=1}^{N}|o_{i}|}\sum_{i=1}^{N}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\gamma_{t}(\theta),\text{clip}\left(\gamma_{t}(\theta),1-\epsilon,1+\epsilon\right)\right]\hat{A}_{i}\right\}\end{split},$ (1)
where
$\gamma_{t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta^{\prime}}(o_{i,t}|q,o_{i,<t})},$ (2)
$\hat{A_{i}}=\frac{r_{i}-\text{mean}(R)}{\text{std}(R)},$ (3)
and $\epsilon$ is clipping hyperparameter to prevent training collapse.
However, current static RL training paradigm still exhibits limitations. Most of the collected rollouts have advantages sharply concentrate near zero, providing limited signals for gradient update. The proportion of rollouts with non-zero gradient continues to decline throughout the training, introducing invalid updates. These limitations indicate the need for a more efficient training framework.
4 Method
In this section, we begin by further analyzing the existing inefficient issues in training. Then, as illustrated in Fig. 3, we introduce Shuffle-R1, which optimizes the training process through two crucial modules: (1) Pairwise Trajectory Sampling and (2) Advantage-based Batch Shuffle.
4.1 Problem Analysis
Advantage Collapsing.
Our probe analysis reveals that, contrary to the ideal scenario, most rollouts exhibit advantages sharply concentrated around zero in current RL paradigm, leading to the Advantage Collapsing phenomenon. Such distribution results in weak gradient signals and hampers effective policy updates, as only rollouts with high-magnitude advantages can provide significant gradient signals. Simply increasing the number of rollouts can collect more amount of valuable rollouts and improve performance to some extent (Fig. 2(a)), but also substantially increase the computational overhead without fundamentally resolving the Advantage Collapsing. These findings highlight the need for a dynamic mechanism that can adaptively identify and select valuable rollouts to improve training efficiency.
Rollout Silencing.
We observe that the proportion of rollouts contributing non-zero gradients steadily decreases as training progresses, resulting in the Rollout Silencing phenomenon. This is primarily caused by the accumulation of factors such as zero advantages, gradient clipping, and overlong truncation, which further exposes the limitations of static sampling paradigm. By tracking queries of varying difficulty, we find that simple queries tend to converge early, while difficult ones consistently yield low accuracy (Fig. 2(b)). Both cases fail to generate diverse and informative rollouts throughout training, contributing to Rollout Silencing. Moreover, in standard training pipelines, each rollout is used only once for a single gradient update, limiting the model to fully exploit valuable rollouts. Consequently, a dynamic sampling strategy that can adaptively discard ineffective rollouts and queries, while allowing repeated learning from valuable rollouts is of great importance.
Figure 2: (a) Model accuracy improves with larger rollout sizes. (b) Queries with different difficulties demonstrate varying accuracy during training, their corresponding rollouts have different diversity and qualities consequently.
Figure 3: Overview of our proposed Shuffle-R1. After advantage calculation, we first conduct Pairwise Trajectory Sampling to obtain valuable trajectory pairs from original rollout pool, then perform Advantage-based Batch Shuffle to reshape the distribution of valid trajectories in a batch.
4.2 Pairwise Trajectory Sampling
To relieve Advantage Collapsing, we seek to select trajectories that offer stronger learning signals. Since increasing the rollout number enlarges the chance of capturing high-advantage samples, we propose Pairwise Trajectory Sampling (PTS), a data-centric module to selectively amplify valuable learning signals. Rather than evaluating trajectories in isolation, PTS organizes candidate rollouts into structured contrastive pairs. This pairing mechanism captures both high and low advantage signals jointly, forming informative “positive-negative” pairs. Only pairs with the largest advantage contrast are then retained for training. This process ensures that limited update bandwidth is focused on trajectories that are both diverse and gradient-rich.
Given a query $q$ and a rollout size of $2N$, the rollout trajectories group is denoted as $O=\{o_{i}\}_{i=1}^{2N}$. The corresponding reward and advantage sets are $R=\{r_{i}\}_{i=1}^{2N}$ and $A=\{\hat{A}_{i}\}_{i=1}^{2N}$, respectively. To identify informative trajectory pairs, our proposed pairing mechanism follows a simple ‘max-and-min’ principle by matching the trajectory with the highest advantage to that with the lowest, the second highest to the second lowest, and so on. We denote the sorted advantage values in descending order as:
$A_{s}=\{\hat{A}_{(i)}\}_{i=1}^{2N},\quad\text{where }\hat{A}_{(1)}\geq\hat{A}_{(2)}\geq\cdots\geq\hat{A}_{(2N)}.$ (4)
Based on this ordering, we construct the pairing set as:
$P=\{(o_{(i)},o_{(2N-i+1)})\}_{i=1}^{N}.$ (5)
In this scheme, the original $2N$ rollouts are easily sorted and reorganized into $N$ pairs. The top-ranked pairs typically consist of trajectories with high-magnitude but opposite-sign advantages, forming contrastive pairs akin to “positive-negative” samples. In contrast, the bottom-ranked pairs involve trajectories with advantages closer to zero.
As implied by Eq. 1, trajectories with higher absolute advantages contribute more significantly to the gradient update, while those with near-zero advantages have negligible impact. We apply a simple top-$k$ sampling strategy to select a subset of valid pairs:
$P_{v}=\{(o_{(i)},o_{(2N-i+1)})\}_{i=1}^{M},\quad M=\alpha N,\ \alpha\in(0,1).$ (6)
where $\alpha$ is a hyperparameter controlling the sampling ratio from the pairing set. We only keep the trajectories in the valid set for gradient update.
By introducing a structured contrastive sampling scheme, PTS enables more effective trajectory selection from a broader exploration space without increasing the gradient computation cost. The contrastive pairing not only filters out low-signal trajectories, but also sharpen the model’s policy gradient through direct comparison. PTS shifts the focus of RL fine-tuning from uniform exploration to gradient-informed selection, representing a principled step toward more efficient data usage in RL training.
4.3 Advantage-based Batch Shuffle
While PTS could mitigate Advantage Collapsing and improve RL training performance, the Rollout Silencing issue remains unresolved. To overcome this issue, we propose Advantage-based Batch Shuffle (ABS) module that dynamically reshapes training batches to prioritize and reinforce high-value samples. Rather than relying on static data flow, ABS adaptively redistributes trajectories within each training batch, enabling more frequent updates to trajectories with high learning utility. Built on top of PTS, it serves to magnify the gradient exposure of informative samples, reshaping the training data distribution to achieve better data utilization and training efficiency.
Denote a data batch:
$B=\{p_{i}^{g}=(o_{i,1}^{g},\hat{A}_{i,1}^{g},o_{i,2}^{g},\hat{A}_{i,2}^{g},q^{g})\}_{i=1\sim M,g=1\sim G},$ (7)
with batch size of $M\times G$, In the standard gradient update process, $B$ is sequentially divided into $K$ mini-batches, each contains $MG/K$ samples.
In our ABS module, we first assign an importance weight to each trajectory pair based on the sum of the absolute advantages:
$W(p_{j})=|\hat{A}_{j,1}|+|\hat{A}_{j,2}|.$ (8)
We then define the sampling probability of each pair $p_{j}\in P_{v}$ as:
$\Phi(p_{j})=\frac{W(p_{j})}{{\sum_{k=1}^{|P_{v}|}}W(p_{k})}.$ (9)
Based on the sampling probability, we perform $S$ sub-sampling from original batch $B$, each sub-sampling has a capacity of $T$ pairs ($2T$ trajectories):
$B_{s}=\{p_{s,t}\}_{t=1}^{T},\quad\text{s.t. }p_{s,t}\neq p_{s,t^{\prime}},\forall\,t\neq t^{\prime}.$ (10)
All the sub-sampling batches are sequentially combined to form the reshuffled batch $B^{\prime}=\bigcup_{s=1}^{S}B_{s}$. During the ABS process, we set $|B^{\prime}|=|B|,\ \text{i.e., }S\times T=MG$ to ensure the reshuffled batch matches the same size as the original batch. The reshuffled batch will maintain the gradient update paradigm of the original method.
The ABS module optimizes the learning process through Advantage-aware shuffling and Inter-sub-batch resampling. It increases the update frequency of trajectories with higher advantages, maintains diversity while reinforcing high-value samples through repeated exposure. Together, these designs transform each batch into a soft-prioritized structure that better reflects training signal utility.
5 Experiments
5.1 Experimental Setup
Datasets and Benchmarks.
We first conduct our experiments on Geometry3K dataset (Lu et al. 2021) ($2.1k$ training samples.) and a subset of MMK12 dataset (Meng et al. 2025) containing the same amount of data, to investigate model performance on limited training resources. Then, we conduct experiments on larger data scale with MM-Eureka dataset (Meng et al. 2025), seeking for better model performance. We randomly select $27k$ samples from MM-Eureka dataset and mix them with Geometry3K dataset, forming a total of $30k$ training corpora. All training samples are in free-form format.
We first perform evaluation on in-domain test set of Geometry3K and MMK12. Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al. 2024b), MathVision (Wang et al. 2024), WeMath (Qiao et al. 2024), MathVista (Lu et al. 2023), HallusionBench (Guan et al. 2024) and ChartQA (Masry et al. 2022). These benchmarks span across math reasoning, visual perception, and chart understanding. We use MathRuler to evaluate questions with free-form ground truths and Gemini-2.0-Flash-001 (Deepmind 2025) to evaluate questions with multi-choice ground truths.
Method Geo3K Math Avg. HallBench ChartQA
Qwen-3B 25.79 41.71 59.83 73.08
+ GRPO 42.64 46.74 63.09 76.20
+ DAPO 45.09 48.08 63.24 76.70
+ Ours 47.88(+22.09) 48.70(+6.99) 63.19(+3.36) 77.04(+3.06)
Qwen-7B 38.12 49.82 65.19 79.84
+ GRPO 52.60 53.13 68.56 80.84
+ DAPO 54.43 54.19 69.29 81.20
+ Ours 55.89(+17.77) 54.63(+4.81) 69.51(+4.32) 81.64(+1.80)
Table 1: Performance of Shuffle-R1 on Geometry3K dataset compared with GRPO and DAPO.
Method K12 Math Avg. HallBench ChartQA
Qwen-3B 42.42 41.71 59.83 73.08
+ GRPO 59.19 48.71 64.14 77.12
+ DAPO 61.42 49.75 65.08 77.00
+ Ours 62.22(+19.80) 50.05(+8.34) 65.72(+5.89) 78.28(+5.20)
Qwen-7B 52.13 49.82 65.19 79.84
+ GRPO 66.15 54.47 67.75 82.48
+ DAPO 68.35 54.52 68.66 82.52
+ Ours 68.78(+16.65) 55.02(+5.20) 69.87(+4.68) 82.60(+2.76)
Table 2: Performance of Shuffle-R1 on K12 dataset compared with GRPO and DAPO.
Implementation Details.
We use EasyR1 (Yaowei et al. 2025) as our training codebase. We choose Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct (Bai et al. 2025) as base model to verify our method’s generalizability on model scales. Parameters of vision encoder are kept frozen. We set update batch size to 128 and rollout batch size ($G$) to 512. Rollout temperature is set to 1.0 and learning rate is set to 1e61e-6. All experiments are conducted on $8\times$ 80G GPUs. For PTS, we sample 4 trajectory pairs (8 trajectories) from 16 rollouts for each query, striking a balance between training cost and exploration space. For ABS, we set the sub-sampling batch size ($T$) to 256 pairs (512 query-response trajectories) and the shuffle times ($S$) to 8. For evaluation, we set the temperature to 0.5 and report the average pass@1 accuracy of 8 tests to reduce randomness.
5.2 Main Results
Model MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
Close-source
GPT-4o (Achiam et al. 2024) 50.8 30.4 63.8 68.8 55.0 - -
o1 (OpenAI 2024) 57.0 60.3 73.9 - - - -
Gemini-2.0 pro (Deepmind 2025) 67.3 48.1 71.3 - 49.8 - -
Claude-3.7-Sonnet (Anthropic 2025) 52.0 41.3 66.8 72.6 55.4 - -
Open-Source SFT
InternVL-2.5-8B (Chen et al. 2025b) 39.5 17.0 64.5 - 50.1 79.1 -
InternVL-3-8B (Zhu et al. 2025) - 29.3 71.6 - 49.9 86.6 -
Qwen2.5-VL-3B (Bai et al. 2025) 34.8 21.9 58.4 51.7 59.8 73.1 49.9
Qwen2.5-VL-7B (Bai et al. 2025) 42.6 25.8 67.4 63.5 65.2 79.8 57.4
Cold Start + RL
R1-VL-7B (Zhang et al. 2025) 40.1 24.3 62.3 59.8 60.9 76.1 53.9
Vision-R1-7B (Huang et al. 2025) 46.1 - 70.8 - 57.8 83.1 -
R1-OneVision-7B (Yang et al. 2025) 43.0 24.8 61.2 60.6 66.4 77.8 55.2
OpenVLThinker-7B (Deng et al. 2025) 46.4 24.8 69.7 67.2 59.1 78.4 57.6
VLAA-Thinker-7B (Chen et al. 2025a) 48.9 26.3 69.9 67.7 67.5 80.1 60.1
Zero RL
$M$-Eureka-Qwen-7B (Meng et al. 2025) 49.6 27.4 70.6 67.4 66.7 79.0 60.1
MMR1-Math-7B (Leng et al. 2025) 39.2 31.9 71.5 70.7 69.6 82.0 60.8
ThinkLite-VL-7B (Wang et al. 2025b) 45.2 28.0 72.4 69.3 70.2 82.0 61.2
VL-Rethinker-7B (Wang et al. 2025a) 51.7 29.7 72.0 70.1 69.9 79.0 62.1
NoisyRollout-7B-K12 (Liu et al. 2025a) 50.1 28.0 70.9 70.8 70.1 81.4 62.1
Shuffle-R1-Qwen-3B (Ours) 44.2 26.8 70.4 66.5 69.2 79.9 59.5
Shuffle-R1-Qwen-7B (Ours) 53.9 30.0 77.0 72.3 71.0 84.1 64.7
Table 3: Model performance on representative visual reasoning benchmarks. Models marked with ‘’ are evaluated using our own evaluation scripts with vLLM. Vision-R1-7B used WeMath and MathVision as training data, its performance on these benchmarks are omitted. Best performance of RL-only models marked with Bold, second best with underline.
Comparison with Representative Algorithms.
We compare model performance of our method with GRPO (Shao et al. 2024) and DAPO (Yu et al. 2025). As shown in Tab. 1, after training on the Geometry3K dataset, the 3B model trained with our method achieved 47.88% accuracy on in-domain test set, with a 5.2% improvement compared to model trained with GRPO and 2.7% improvement compared to DAPO. Similarly, for the 7B model, our method led to 3.3% improvement over GRPO and 1.4% over DAPO, achieving 55.89% accuracy. The superiority of our framework is further highlighted in out-of-domain benchmarks. For out-of-domain math reasoning task, the average accuracy is improved by 1.96% on 3B model and 1.5% on 7B model compared to GRPO, and surpasses DAPO as well. Similar performance gains are also observed on HallusionBench and ChartQA, where the data gap is even further from training data, showing strong generalizability of our framework. Tab. 2 demonstrates similar performance results on K12 dataset, where consistent performance gains are achieved on both in-domain and out-of-domain benchmarks, highlighting our framework’s generalizability on different data distribution.
Comparison with RL-based models.
We conduct larger scale experiments on MM-Eureka dataset. As shown in Tab. 3, trained with $30k$ selected data from diverse sources for 150 steps, our 7B model exhibits a substantial accuracy gain over the base model (Qwen2.5-VL-7B). Moreover, it outperforms a series of open-source 7B competitors who also adopt RL training strategies, e.g. MM-Eureka with direct RL and VLAA-Thinker with RL after cold-start SFT. Notably, our model achieves competitive or superior performance on several benchmarks compared to leading close-source models, for instance, Claude-3.7-Sonnet (Anthropic 2025) and GPT-4o (Achiam et al. 2024). Under the same setting, our 3B variant also demonstrates strong performance, even outperforming several 7B models on certain benchmarks. These results highlight the superiority of our proposed approach in boosting the training efficiency in reinforcement learning.
Figure 4: Advantage distribution in a training batch of GRPO and our framework.
Efficiency Analysis.
The improvement in model performance mainly stems from better training efficiency. Fig. 2(a) in previous section shows that, under the same update rollout size, model trained under our framework significantly outperforms GRPO, indicating more effective utilization of training data. Advantage distribution analysis in Fig. 4 proves that, PTS effectively mitigates Advantage Collapsing by increasing the proportion of large-magnitude advantages. ABS further optimizes the batch composition, enabling the model to focus on more informative trajectories.
Fig. 5(a) and (b) further probe into training dynamics and demonstrate that our framework consistently achieves higher training and validation accuracy, reaching comparable performance as GRPO but with only half the training steps. Moreover, our framework effectively mitigates the issue of “Rollout Silencing” shown in Fig. 5(c), maintaining a high token utilization rate across all training stages. Fig. 5(d) further illustrates the favorable trade-off between training scale and computational cost of our approach, substantially expanding the RL exploration space with minimal additional overhead.
5.3 Ablation Study
Figure 5: (a): Training accuracy of GRPO and Shuffle-R1. (b): Validation accuracy of GRPO and Shuffle-R1. (c): Token utilization rate of GRPO and Shuffle-R1. (d): Shuffle-R1 achieves better performance with minimal extra time cost.
We conduct ablation experiments on Qwen2.5-VL-3B-Instruct using Geometry3K, focusing on two objectives: (1) assessing the contribution of each component in our framework, and (2) validating the impact of key designs.
GRPO PTS ABS Geo3k Math Avg. HallBench ChartQA
25.79 41.71 59.83 73.08
42.64 46.74 63.09 76.20
46.21 47.64 63.40 76.52
47.88 48.70 63.19 77.04
Table 4: Ablation study on effectiveness of PTS and ABS.
Effectiveness of Algorithm Design.
We assess the effectiveness of proposed PTS and ABS modules by successively introduce them to the baseline. As shown in Tab. 4, in the in-domain Geometry3K test set, PTS demonstrates significant performance gains, improving the model accuracy from 42.64% to 46.21% (+3.57%). When ABS is introduced for co-optimization, the model achieves a further 1.67% performance gain, eventually reaching an accuracy of 47.88%. This improvement trend continues in the out-of-domain evaluation scenario: in the math reasoning task, the full setting (PTS + ABS) achieves an absolute performance improvement of 2.0% compared to GRPO (48.70% vs. 46.74%, with 41.71% before post-training). In the ChartQA dataset, which has more significant distributional differences, the joint application of PTS and ABS still maintains an effective performance gain (77.04% vs. 76.20%, with 73.08% before post-training). Our method also demonstrates improved performance in HallusionBench.
Rationality of Pairwise Trajectory Sampling.
One core mechanism of the PTS lies in the structured contrastive sampling scheme. To verify the effectiveness of this bidirectional optimization mechanism, we designed the following experiments: (1) one-way positive sampling (only select trajectories with highest advantage); (2) one-way negative sampling (only select trajectories with lowest advantage); and (3) unbiased random sampling. All settings maintained a consistent sampling ratio (8 valid trajectories from 16 rollouts) to ensure fairness. We disable ABS since it relies on the pairing result of PTS. As shown in Tab. 5, model trained with PTS receives a consistent performance gain on both in-domain and out-of-domain tasks, while one-way positive and one-way negative sampling show a decline in all tasks, demonstrating the effectiveness and rationality of our design. Unbiased random sampling only receives minor improvement over baseline, far behind the effectiveness of PTS.
Setting Geo3k Math Avg. HallBench ChartQA
Qwen2.5-VL-3B 25.79 41.71 59.83 73.08
+ GRPO 42.64 46.74 63.09 76.20
Ablation on PTS
+ only max 41.26 44.77 63.30 75.64
+ only min 23.36 41.52 60.98 74.36
+ rand. select 43.53 46.62 63.19 76.00
+ PTS 46.21 47.64 63.40 76.52
Ablation on ABS
+ rand. shuffle 46.05 47.40 63.19 76.60
+ reorder 46.28 47.64 63.09 76.64
+ ABS 47.88 48.80 63.19 77.04
Table 5: Ablation study on rationality of PTS and ABS.
Rationality of Advantage-based Batch Shuffle.
ABS introduces Advantage-aware shuffling and Inter-sub-batch resampling to reshape the training batch. To validate their effectiveness, we designed two contrastive experiments: (1) unbiased shuffle: using uniformly distributed sampling weights to perform shuffle strategy. and (2) static reorder: randomly reorder the training batch without shuffle, maintaining the original data distribution. We enable PTS during training as ABS relies on its pairing result. As shown in Tab. 5, model trained with ABS significantly outperforms the contrastive settings. The unbiased shuffle setting even performs worse compared to the PTS-only setting, demonstrating the rationality of ABS. The static reorder setting has no improvement compared to PTS-only setting, as they have the same data distribution.
Hyperparameters.
We investigate the impact of two key hyperparameters in our framework: (1) the sampling ratio ($\alpha$) in PTS, and (2) the shuffle times ($S$) in ABS. For $\alpha$, we fix $S=8$ and test values of 0.25, 0.5, and 0.75 (i.e., selecting 4, 8, and 12 samples from 16 rollouts). As shown in Fig. 6(a), both $\alpha=0.75$ and $\alpha=0.5$ yield strong performance, while $\alpha=0.25$ lags behind. We attribute this to over-pruning, where filtering out rollouts too aggressively may reduce data diversity. We choose $\alpha=0.5$ to strike a balance between signal quality and computational efficiency. For $S$, we fix $\alpha=0.5$ and vary the shuffle times as 4, 8, and 16. Fig. 6(b) shows that performance improves with increasing $S$, but saturates beyond $S=8$. This suggests that moderate resampling enhances data exposure, but too many shuffles may offer diminishing returns.
Figure 6: Ablation on key hyper parameters. (a): Effect of sampling ratio $\alpha$. (b): Effect of shuffle times $S$.
6 Conclusion
In this paper, we propose Shuffle-R1, a simple but effective framework that improves the training efficiency of reinforcement learning of multimodal large language models. Through Pairwise Trajectory Sampling and Advantage-based Batch Shuffle, our framework significantly outperforms representative algorithms and models in both in-domain and out-of-domain tasks, demonstrating the value of data-centric adaptive design. We hope that our motivations, method, and findings are helpful for further research.
Appendix A Pseudo Code
Here, we provide the pseudo code of Shuffle-R1 for readers to better understand the pipeline flow, and for better transparency in algorithm understanding and reproducibility.
Algorithm 1 Shuffle-R1 Workflow
1:Input: $\mathcal{Q}$ (queries), $\pi_{\theta}$ (policy), $2N$ (rollouts per query), $\alpha$ (sampling ratio), $S$ (shuffle rounds)
2:Output: Optimized batch $\mathcal{B}^{\prime}$ for gradient update
3:Initialize global batch $\mathcal{B}\leftarrow\emptyset$
4:for each $q\in\mathcal{Q}$ do $\triangleright$ PTS Phase
5:  Generate $\{o_{i}\}_{i=1}^{2N}\sim\pi_{\theta}(q)$
6:  Compute $\{\hat{A}_{i}\}$ via $R=RewardFunc(\{o_{i}\},q)$
7:  Sort pairs: $\{(o_{(i)},o_{(2N-i+1)})\}_{i=1}^{N}\leftarrow\text{MaxMinPair}(\{\hat{A}_{i}\})$
8:  Retain top-$\lfloor\alpha N\rfloor$ pairs: $\mathcal{B}_{q}\leftarrow\{(o_{(k)},o_{(2N-k+1)})\}_{k=1}^{\lfloor\alpha N\rfloor}$
9:  Aggregate: $\mathcal{B}\leftarrow\mathcal{B}\cup\mathcal{B}_{q}$
10:end for
11:Compute $W_{j}=|\hat{A}_{p}|+|\hat{A}_{q}|$ $\forall(o_{p},o_{q})\in\mathcal{B}$ $\triangleright$ ABS Phase
12:Calculate $P_{j}={W_{j}}/{\sum W_{k}}$ for weighted sampling
[…]
14:return $\mathcal{B}^{\prime}$ $\triangleright$ Jointly optimized training data
Appendix B Prompt Design
[…]
Appendix C Experiment Settings
We report details of our training and evaluation settings here, including reward function design, main hyperparameters and computing resources.
Reward Calculation.
We adopt a combination of format reward and accuracy reward as the final reward in reinforcement learning. The format reward and accuracy reward are calculated as follows:
$r_{\text{format}}=\begin{cases}1,&\text{if format is {correct}}\\ 0,&\text{if format is {incorrect}}\end{cases}$ (11)
[…]
The final reward is the weighted sum of above rewards:
$r_{\text{overall}}=0.1\times r_{\text{format}}+0.9\times r_{\text{acc}}$ (13)
Format reward is assigned to a smaller weight since response formatting is easy to learn.
Hyperparameters.
We use EasyR1 (Yaowei et al. 2025) as our training framework. Full hyperparameter settings during training is shown in Tab. 6. For experiments on Geometry3K and K12 with $\sim 2.1k$ training samples, we set the training steps to 80. For the joint training experiments ($\sim 30k$ training samples), we increase the training steps to 150 due to extended data size. Other hyperparameters that are not mentioned are kept to default values of EasyR1.
Hyperparameters Value
max pixels 1000000
min pixels 262144
max prompt length 2048
max response length 2048
rollout batch size 512
global batch size 128
learning rate 1e-6
optimizer AdamW
rollout temperature 1.0
rollout top p 0.99
evaluation temperature 0.5
rollout group number 16
PTS sampling ratio ($\alpha$) 0.5
ABS shuffle times ($S$) 8
KL coefficient 0.0
vision encoder frozen
Table 6: Hyperparameter settings.
Computing Resources.
All experiments are conducted on $8\times$ 80G-SXM GPUs.
Appendix D Dataset Collection
For training dataset, we choose Geometry3K (Lu et al. 2021) and MMK12 dataset. Geometry3K dataset is a high quality real-world geometry problem solving dataset, containing 2.1K training samples and 601 test samples. It contains a wide range of geometry problems with varying levels of difficulty. The text problems are very compact with most of the geometric conditions represented in images, making it suitable for our training. MMK12 dataset is introduced by MM-Eureka (Meng et al. 2025), which contains 16k math reasoning training samples. The training samples have both geometric and non-geometric problems and have been carefully examined and manually filtered to ensure quality. The test set of MMK12 further includes other STEM problems such as physics, biology and chemistry. During our training, we used a randomly selected subset of MMK12 provided by NoisyRollout (Liu et al. 2025a) (referred to as ‘K12’ in the paper to distinguish it from full set of MMK12), which has 2.1k samples with the same size as Geometry3K. All the training samples are in free-form format. Examples of training data shown in Fig. 7.
Figure 7: Top: Examples of Geometry3K. Bottom: Examples of MMK12.
To fully evaluate model’s reasoning ability, apart from in-domain test set from training dataset, we select several representative benchmarks to examine model’s performance on math reasoning, visual perception, and chart understanding tasks. For math reasoning task, we choose MathVerse (Zhang et al. 2024b), MathVision (Wang et al. 2024), WeMath (Qiao et al. 2024) and MathVista (Lu et al. 2023). We choose HallusionBench (Guan et al. 2024) and ChartQA (Masry et al. 2022) for visual perception and chart understanding tasks respectively. We believe that a good RL algorithm can not only improve in-domain performance, but also have the potential to promote robust generalization to out-of-domain tasks. Examples of evaluation data shown in Fig. 8.
Figure 8: Top: Examples of math reasoning tasks. Middle: Examples of visual perception tasks. Bottom: Examples of chart understanding tasks.
Appendix E Evaluation Metrics
We adopt pass@1 accuracy as evaluation metric. During evaluation, we set the decoding temperature to 0.5 and perform 8 independent runs, reporting the average pass@1 accuracy as final metric. The choice of temperature doesn’t degrade model performance, and the averaged accuracy reduces the randomness, resulting in a more stable and reliable evaluation. Prompt format for evaluation is kept identical to training prompt. Evaluation settings of our model and all the reproduced results in main paper are kept the same to ensure a fair comparison.
For evaluation of MathVerse, MathVision, WeMath, MathVista and ChartQA, we employ Gemini-2.0-Flash-001 (Deepmind 2025) to first extract the predicted answer from model response then compare it with ground truth. Fig. 10 and Fig. 10 demonstrate the extraction and verification prompt for Gemini. Specifically, we report accuracy on WeMath under loose mode, and overall accuracy on MathVerse (including all sub categories: Text Dominant, Text Lite, Vision Intensive, Vision Dominant and Vision Only).
Figure 9: Extraction prompt for Gemini.
Figure 10: Score prompt for Gemini.
Appendix F More Experimental Results
Detailed Model Performance.
We provide a more detailed performance of models trained on Geometry3K and K12 dataset in Tab. 7, reporting model performance on each out-of-domain benchmarks as a supplement to “Math Avg.” columns in main paper. Trained with only $2.1k$ data, both the 3B and 7B model demonstrate significant performance gains.
Comparison with more RL algorithms.
We provide additional comparisons with more representative RL algorithms, i.e. RLOO (Ahmadian et al. 2024) and REINFORCE++ (Hu 2025). We conduct experiments on Qwen2.5-VL-3B-Instruct with Geometry3K dataset. All the experiment settings are kept the same as GRPO and DAPO in main paper. The final model performance is shown in Tab. 8. Our framework outperform these algorithms by a large margin in both in-domain and out-of-domain tasks.
Figure 11: Performance of Shuffle-R1 compared with directly applying Prioritized Experience Replay.
Method MathVerse MathVision MathVista WeMath HallBench ChartQA Total Avg.
Qwen2.5-VL-3B 34.77 21.94 58.40 51.72 59.83 73.08 49.96
+ Ours (Geo3K) 43.55 25.30 61.80 64.14 63.19 77.04 55.84
+ Ours (K12) 44.06 26.48 64.90 64.77 65.72 78.28 57.37
Qwen2.5-VL-7B 42.59 25.76 67.40 63.51 65.19 79.84 57.38
+ Ours (Geo3K) 50.96 27.47 70.90 69.19 69.51 81.64 61.61
+ Ours (K12) 48.59 28.61 73.20 69.71 69.87 82.60 62.09
Table 7: Detailed performance on out-of-domain benchmarks of models trained on Geometry3K and K12 data. Highest accuracy marked in Bold.
Method Geo3K MathVerse MathVision MathVista WeMath HallBench ChartQA
Qwen2.5-VL-3B 25.79 34.77 21.94 58.40 51.72 59.83 73.08
+ RLOO 42.09 39.94 22.96 58.90 59.48 64.24 76.68
+ REINFORCE++ 41.76 41.70 24.86 60.90 63.51 62.99 76.20
+ Ours 47.88 43.55 25.30 61.80 64.14 63.19 77.04
Table 8: Performance of Shuffle-R1 on Geometry3K dataset compared with RLOO and REINFORCE++. Highest accuracy marked in Bold.
Comparison with Direct Experience Replay.
We conduct a comparative study by replacing ABS with a prioritized experience replay mechanism. Experience replay maintains a decoupled buffer of past samples, whereas ABS adopts an online, in-place shuffle strategy to dynamically reshape the data distribution. As shown in Fig. 11(a) and (b), in the later stages, the experience replay setting exhibits a plateau in training accuracy and even a drop in validation accuracy, indicating potential overfitting to stale samples. This suggests that prioritized experience replay may overly emphasize historical trajectories, leading to suboptimal convergence. Moreover, ABS proves to be more effective in mitigating the Rollout Silencing.
Appendix G Case Study
Here we provide some qualitative case study of Shuffle-R1’s reasoning outputs.
Fig. 13 and Fig. 13 show two improved cases on math related tasks. In case 1, the base model has an error in visual information parsing, referring to angles not existed in the figure, resulting in a reasoning fault. The RL model correctly parsed and solved the geometry problem. In case 2, the base model misused a geometric theorem, leading to wrong answer, while the RL model correctly solved the problem with accurate theorem.
Fig. 15 demonstrates improved reasoning ability in RL model, where the base model has accurate perception about the chart but has reasoning error in CoT. The RL model can not only accurately parse the visual information, but also perform correct reasoning to finally reach the right answer.
Case 4 in Fig. 15 further demonstrates that RL model also has better visual perception ability compared to base model. The figure from HallusionBench has been artificially inserted with an image of a hen, which only occupies a very small region of the original image. This modification has resulted in a perception error and in base model, but the RL model can accurately identify the inserted image.
Figure 12: A case study on math related task.
Figure 13: Another case study on math related task.
Figure 14: A case study on chart understanding task.
Figure 15: A case study on visual perception task.

### v4

@@
 It introduces (1) Pairwise Trajectory Sampling, which selects high-contrast trajectories with large advantages to improve gradient signal quality, and (2) Advantage-based Batch Shuffle, which increases exposure of valuable rollouts through strategic batch reshuffling.
-Experiments across multiple reasoning benchmarks demonstrate that our framework consistently outperforms strong RL baselines with minimal computational overhead.
-These results highlight the importance of data-centric adaptations for more efficient RL training in MLLM.
-footnotetext: 🖂 Corresponding author.
-   This work was done during the research internship of Linghao Zhu, Yiran Guan, and Dingkang Liang at Xiaomi.
+Experiments across multiple reasoning benchmarks show that our framework consistently outperforms strong RL baselines with minimal computational overhead.
+These results support the potential of data-centric adaptations for more efficient RL training for MLLM.
 1 Introduction
-Inspired by the slow and deliberate thinking process in human cognition, reasoning capabilities empower Large Language Model (LLM) to plan, reflect, and generalize beyond mere memorization (Chu et al.
-2025a; Kang et al.
-2024).
-Recent advances have demonstrated that integrating reinforcement learning (RL) into the training of LLMs can significantly enhance such capabilities, particularly in complex domains such as mathematical problem solving and code generation (Guo et al.
-2025; Deepmind 2025; OpenAI 2024; Seed et al.
-2025).
-Notably, DeepSeek-R1 (Guo et al.
-2025) leverage reward signals derived exclusively from verifiable outcomes to yield impressive performance gains.
-Beyond textual tasks, RL has also seen increasing application in multimodal domains, including object detection, segmentation, and video understanding (Li et al.
-2025; Liu et al.
-2025c, b; Wang et al.
-2025c), highlighting its potential to support generalizable reasoning across modalities.
-Figure 1: (a): Advantage Collapsing, where most advantages concentrate near zero.
-(b): Rollout Silencing, where the ratio of rollouts with non-zero gradient consistently drops.
-The phenomenon gets worse in larger models.
-Despite the growing interest in RL for LLM and Multimodal LLM (MLLM), several practical challenges remain unresolved, including slow convergence, training instability, and suboptimal efficiency.
-Recent studies have proposed various improvements, such as empirical optimization of training configurations (Yu et al.
-2025), integration of experience replay (Wang et al.
-2025a), and refinement of objective functions (Chu et al.
-2025b).
-Other works focus on reward optimization (Liu et al.
-2025a; Ma et al.
-2025).
-However, most of these approaches remain confined to the standard RL training paradigm, where trajectories are uniformly sampled without considering their varying informativeness or difficulty.
-This static strategy overlooks a crucial insight that not all learning signals are equally valuable, which leads us to consider a fundamental question: Can dynamically prioritizing trajectories provide richer gradient information and lead to more efficient training?
-Motivated by this question, we examine the current RL training practices and identify two critical yet might underexplored inefficiencies.
-First, Advantage Collapsing emerges when most computed advantages cluster excessively near zero, drowning out informative signals from trajectories with large-magnitude advantages, resulting in weak or negligible gradient updates (Fig.
-1(a)).
+Reinforcement learning (RL) has emerged as a powerful tool to enhance large language models (LLMs) to plan, reflect, and generalize, enabling stronger performance in complex reasoning domains such as mathematical problem solving and code generation (Guo et al., 2025; Deepmind, 2025; OpenAI, 2024; Seed, 2025).
+Notably, DeepSeek-R1 (Guo et al., 2025) leverages reward signals derived exclusively from verifiable outcomes to yield impressive performance gains.
+Beyond textual tasks, RL has also seen increasing application in various multimodal domains (Li et al., 2025; Liu et al., 2025c; b; Wang et al., 2025c), highlighting its potential to support generalizable reasoning across modalities.
+To better incorporate RL in LLM and Multimodal LLM (MLLM), recent studies have proposed various improvements, including better framework optimization (Yu et al., 2025; Wang et al., 2025a; Chu et al., 2025b) and more sophisticated reward optimization (Liu et al., 2025a; Ma et al., 2025).
+However, most approaches remain confined to static sampling paradigm, where trajectories are sampled and treated uniformly.
+Such strategy overlooks a crucial insight that not all learning signals are created equal and their informativeness varies and evolves during training.
+Ignoring variation in signal quality risks the training process being overwhelmed by noisy trajectories and underusing the truly useful signals.
+An ideal framework should instead concentrate updates on golden signals while filtering out noisy ones.
+Consequently, it leads us to consider a fundamental question: Can dynamically prioritizing trajectories provide richer gradient information and lead to more effective training?
+Motivated by this question, we investigate into current RL training practices and reveal two critical yet might underexplored limitations.
+First, Advantage Collapsing emerges when most computed advantages cluster excessively near zero, drowning out informative signals from trajectories with large-magnitude advantages, resulting in extremely weak or negligible gradient updates and wiping out decisive ones (Fig.
+2(a)).
 Second, Rollout Silencing arises as the fraction of rollouts contributing non-zero gradients steadily declines during training (Fig.
-1(b)), leading to wasted computation and underutilization of data.
+2(b)), leading to catastrophic waste of computation without fully utilizing informative signals.
 These two findings demonstrate a pressing need for adaptive mechanisms to prioritize, reuse, and reallocate gradient exposure toward informative samples.
-In this paper, we introduce Shuffle-R1, an innovative framework designed to dynamically prioritize and amplify critical gradient information.
-Rather than asking how to design better rewards, we try to ask what data should the model actually update on during RL fine-tuning.
-Specifically, Shuffle-R1 introduces two elegantly simple yet highly effective modules: (1) Pairwise Trajectory Sampling, which selects high-contrast trajectory pairs with large-magnitude advantages from an extended rollout pool, concentrating learning signals to mitigate Advantage Collapsing; and (2) Advantage-based Batch Shuffle, which dynamically reshapes training batches to prioritize informative trajectories while down-weighting ineffective ones, alleviating Rollout Silencing to improve data utilization.
-Our framework is lightweight, modular, and seamlessly integrable with existing RL algorithms.
-Experiments demonstrate that Shuffle-R1 significantly improves model performance across challenging multimodal reasoning tasks, even surpassing GPT-4o and Claude-3.7 on MathVerse and MathVista.
-Moreover, it achieves competitive performance to GRPO while requiring only half of the training steps.
-In summary, our contributions are three-fold: 1) We reveal two critical yet underexplored limitations that undermine training efficiency in RL fine-tuning for MLLM, i.e., Advantage Collapsing and Rollout Silencing.
-2) We propose Shuffle-R1, a novel and adaptive RL framework that dynamically selects high-contrast trajectories and reshapes training batches to emphasize informative samples.
-3) Extensive experiments across model scales and both in-domain and out-of-domain benchmarks demonstrate the effectiveness and generalizability of our framework.
+Figure 1: (a) Advantage Collapsing, where most advantages concentrate near zero. (b) Rollout Silencing, where the ratio of rollouts with non-zero gradient consistently drops.
+Figure 2: Pipeline comparison. (a) Static paradigm. (b) Rule-based pre-filter paradigm. (c) Dynamic paradigm can ‘interact’ with model during training.
+In this paper, we present Shuffle-R1, a framework that dynamically prioritizes and amplifies critical gradient signals during RL fine-tuning.
+Guided by the philosophy that what data the model updates on is as important as how it updates, Shuffle-R1 introduces two effective modules: (1) Pairwise Trajectory Sampling, which selects high-contrast trajectory pairs with large advantages gaps from an extended rollout pool, concentrating on the most discriminative learning signals to mitigate Advantage Collapsing; and (2) Advantage-based Batch Shuffle, which adaptively reshapes training batches to emphasize informative trajectories while down-weighting unhelpful ones, alleviating Rollout Silencing by removing noisy signals and improving computation utilization.
+Together, our framework embodies the principle of dynamic data prioritization (Fig.
+2), enabling adaptive interaction between model and data.
+Experiments show that Shuffle-R1 substantially improves multimodal reasoning performance, surpassing GPT-4o (Achiam et al., 2024) and Claude-3.7 (Anthropic, 2025) on MathVerse (Zhang et al., 2024b) and MathVista (Lu et al., 2023), and matches GRPO (Shao et al., 2024) with only half the training steps.
+In summary, our contributions are three-fold:
+  • We reveal two critical yet underexplored limitations that undermine training efficiency in RL fine-tuning for MLLM, i.e., Advantage Collapsing and Rollout Silencing.
+  • We propose Shuffle-R1, a novel and adaptive RL framework that dynamically selects high-contrast trajectories and reshapes training batches to emphasize informative samples.
+  • Extensive experiments across model scales and both in-domain and out-of-domain benchmarks demonstrate the effectiveness and generalizability of our framework.
+These results underscore the importance of rethinking which data to update on in RL post-training, moving beyond reward design toward dynamic and adaptive data structuring for more effective reasoning enhancement.
 2 Related Work
@@
 Researchers have explored various approaches to equip LLM with reasoning ability.
-Some early studies performed SFT on complex long chain-of-thought data, leading to performance gains on reasoning tasks (Muennighoff et al.
-2025; Ye et al.
-2025; Yue et al.
-2024; Guo et al.
-2024).
-However, some claim that SFT merely enables the model to memorize the format of reasoning steps and long chains of thought, without fully grasping the ability to reason independently (Chu et al.
-2025a; Kang et al.
-2024; Ye et al.
-2024; Allen-Zhu and Li 2024).
-Some researchers control the model to generate structured chain-of-thought instead of free generation, achieving systematic step-by-step reasoning output (Xu et al.
-2025; Wu et al.
-2025; Thawakar et al.
-2025) Other works attempted to use test-time scaling like Monte Carlo Tree Search (MCTS) (Yao et al.
-2023; Zhang et al.
-2024a; Du et al.
-2024; Yao et al.
-2024) to facilitate complex reasoning by actively extending the output of the model.
-Recently, models such as OpenAI o1/o3 (OpenAI 2024), DeepSeek-R1 (Guo et al.
-2025), Seed-Thinking (Seed et al.
-2025), and Kimi-k1.5 (Team et al.
-2025) utilized RL to enable the model to explore independently, stimulating reasoning ability.
-In particular, DeepSeek-R1-Zero directly conducted RL on pre-trained model without instruction fine-tuning, with verifiable outcome reward functions to replace reward models, achieving surprising reasoning ability.
-The training algorithms for RL are also constantly being optimized (Yu et al.
-2025; Chu et al.
-2025b).
-Our work focuses on a deeper investigation of the efficiency of RL training and puts forward an effective solution to improve both the efficiency and performance of RL training.
+Some early studies performed SFT on complex long chain-of-thought data, leading to performance gains on reasoning tasks (Muennighoff et al., 2025; Ye et al., 2025; Guo et al., 2024).
+However, it has been argued that SFT merely enables the model to memorize the format of reasoning steps and long chains of thought, without fully grasping the ability to reason independently (Chu et al., 2025a; Kang et al., 2024).
+Some researchers control the model to generate structured chain-of-thought instead of free generation, achieving systematic step-by-step reasoning output (Xu et al., 2025; Wu et al., 2025; Thawakar et al., 2025) Other works attempted to use test-time scaling like Monte Carlo Tree Search (MCTS) (Yao et al., 2023; Zhang et al., 2024a; Yao et al., 2024) to facilitate complex reasoning by actively extending the output of the model.
+Recently, models such as OpenAI o1/o3 (OpenAI, 2024), DeepSeek-R1 (Guo et al., 2025), Seed-Thinking (Seed, 2025), and Kimi-k1.5 (Kimi, 2025) utilized RL to enable the model to explore independently, stimulating reasoning ability.
+In particular, DeepSeek-R1-Zero directly conducted RL on pre-trained model without instruction fine-tuning, with verifiable outcome reward functions to replace reward models, achieving remarkable reasoning ability.
+The training algorithms for RL are also constantly being optimized (Yu et al., 2025; Chu et al., 2025b).
+Our work focuses on a deeper investigation of the efficiency of RL training and proposes an effective solution to improve both the efficiency and performance of RL training.
 2.2 Reinforcement Learning for MLLM
-Following the success of DeepSeek-R1, a series of studies have transplanted RL into the training of MLLM and downstream visual tasks, such as Open Vocabulary Object Detection (Liu et al.
-2025c), Reasoning Segmentation (Liu et al.
-2025b), Video Understanding (Li et al.
-2025), Video Localization (Wang et al.
-2025c), etc.
-These works mainly focus on the applicability of RL in downstream tasks.
-Some other works focusing on improving the general reasoning ability of MLLM have achieved performance improvement on reasoning tasks by collecting a large amount of high-quality data (Meng et al.
-2025; Huang et al.
-2025; Yang et al.
-2025; Zhang et al.
-2025; Peng et al.
-2025; Tan et al.
-2025).
+Following the success of DeepSeek-R1, a series of studies have transplanted RL into the training of MLLM and downstream visual tasks, such as Open Vocabulary Object Detection (Liu et al., 2025c), Reasoning Segmentation (Liu et al., 2025b), Video Understanding (Li et al., 2025), Video Localization (Wang et al., 2025c), etc.
+These works mainly focus on the applicability of RL to downstream tasks.
+Some other works focusing on improving the general reasoning ability of MLLM have achieved performance improvements on reasoning tasks by collecting a large amount of high-quality data (Meng et al., 2025; Huang et al., 2025; Yang et al., 2025; Zhang et al., 2025; Peng et al., 2025; Tan et al., 2025).
 These works mainly focus on the organization of high-quality reasoning data and the balance between SFT and RL in the training process.
-Some researchers who conduct in-depth research on the RL mechanism have optimized the RL training process from various aspects, including adding contrastive reward mechanism (Li et al.
-2025), actively introducing reflection tokens during rollouts (Wang et al.
-2025a), optimizing the RL objective function and gradient update mechanism (Chu et al.
-2025b), introducing more diverse rollouts (Liu et al.
-2025a; Yao et al.
-2025), etc.
+Some researchers who conduct in-depth research on the RL mechanism have optimized the RL training process from various aspects, including adding contrastive reward mechanism (Li et al., 2025), actively introducing reflection tokens during rollouts (Wang et al., 2025a), optimizing the RL objective function and gradient update mechanisms (Chu et al., 2025b), and introducing more diverse rollouts (Liu et al., 2025a; Yao et al., 2025).
 The core objective of these works is to optimize the RL training process.
-In our work, we propose a novel training framework that introduces dynamic and adaptive selection and resampling to queries and rollouts, reshaping data distribution for better training efficiency and model performance.
-3 Preliminaries
-Reinforcement Learning in LLM.
-Reinforcement learning is originally introduced as a post-training technique to align LLM outputs with human preference.
-Recently, it has been used in fine-tuning stage to improve the reasoning ability of LLMs.
-Policy gradient algorithm is the most commonly used RL method, whose objective function is to maximize the expectation of return the model gets from the environment.
-Given a query $q$, multiple independent sampled responses $O=\{o_{1},o_{2},...,o_{N}\}$ are sampled from the old policy model $o\sim\pi_{\theta^{\prime}}(q)$, with rollout size $N$.
-Subsequently, the corresponding rewards $R=\{r_{1},r_{2},...,r_{N}\}$ are obtained for each response, in our case they are calculated by verifiable reward functions.
-Advantages $\hat{A}=\{A^{\pi_{\theta}}(r_{1}),A^{\pi_{\theta}}(r_{2}),...,A^{\pi_{\theta}}(r_{N})\}$ are calculated to assess the improvement over the current policy model.
-Importance sampling and clipping are introduced to stabilize the training.
-The core objective function is defined as:
-$\begin{split}&\mathcal{J}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta^{\prime}}(\cdot|q)}\\ &\frac{1}{\sum_{i=1}^{N}|o_{i}|}\sum_{i=1}^{N}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\gamma_{t}(\theta),\text{clip}\left(\gamma_{t}(\theta),1-\epsilon,1+\epsilon\right)\right]\hat{A}_{i}\right\}\end{split},$ (1)
+In our work, we propose a novel training framework that introduces dynamic and adaptive selection and resampling of queries and rollouts, reshaping the data distribution for better training efficiency and model performance.
+3 Method
+In this section, we begin by further analyzing the existing drawbacks in training.
+Then, as illustrated in Fig.
+4, we introduce Shuffle-R1, which optimizes the training process through two crucial modules: (1) Pairwise Trajectory Sampling and (2) Advantage-based Batch Shuffle.
+3.1 Preliminaries
+Policy gradient algorithm is a widely used RL method for MLLM, aiming to maximize the expectation of return from the environment.
+For a given query $q$ , $N$ independent responses $O=\{o_{1},o_{2},...,o_{N}\}$ are sampled from the old policy model $o\sim\pi_{\theta^{\prime}}(q)$.
+Each response receives corresponding reward $R=\{r_{1},r_{2},...,r_{N}\}$ computed via verifiable reward functions.
+Advantages $\hat{A}=\{\hat{A}_{1},\hat{A}_{2},...,\hat{A}_{N}\}$ are then estimated to guide policy updates.
+The core objective is defined as:
+$\mathcal{J}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{N}\sim\pi_{\theta^{\prime}}(\cdot|q)}\frac{1}{\sum_{i=1}^{N}|o_{i}|}\sum_{i=1}^{N}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\gamma_{t}(\theta)\hat{A}_{i},\text{clip}\left(\gamma_{t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right]\right\},$ (1)
 where
-$\gamma_{t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta^{\prime}}(o_{i,t}|q,o_{i,<t})},$ (2)
-$\hat{A_{i}}=\frac{r_{i}-\text{mean}(R)}{\text{std}(R)},$ (3)
-and $\epsilon$ is clipping hyperparameter to prevent training collapse.
-However, current static RL training paradigm still exhibits limitations.
-Most of the collected rollouts have advantages sharply concentrate near zero, providing limited signals for gradient update.
-The proportion of rollouts with non-zero gradient continues to decline throughout the training, introducing invalid updates.
-These limitations indicate the need for a more efficient training framework.
-4 Method
-In this section, we begin by further analyzing the existing inefficient issues in training.
-Then, as illustrated in Fig.
-3, we introduce Shuffle-R1, which optimizes the training process through two crucial modules: (1) Pairwise Trajectory Sampling and (2) Advantage-based Batch Shuffle.
+$\gamma_{t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta^{\prime}}(o_{i,t}|q,o_{i,<t})},\qquad\hat{A_{i}}=\frac{r_{i}-\text{mean}(R)}{\text{std}(R)}$ (2)
+and $\epsilon$ is a clipping hyperparameter to prevent training collapse.
+Despite its practicality, this static RL paradigm has notable drawbacks: advantages often concentrate near zero, yielding weak gradient signals, and the fraction of rollouts with non-trivial updates diminishes as training progresses.
+These issues highlight the need for a more flexible, dynamic and efficient training framework.
 3.2 Problem Analysis
 Advantage Collapsing.
-Our probe analysis reveals that, contrary to the ideal scenario, most rollouts exhibit advantages sharply concentrated around zero in current RL paradigm, leading to the Advantage Collapsing phenomenon.
-Such distribution results in weak gradient signals and hampers effective policy updates, as only rollouts with high-magnitude advantages can provide significant gradient signals.
-Simply increasing the number of rollouts can collect more amount of valuable rollouts and improve performance to some extent (Fig.
-2(a)), but also substantially increase the computational overhead without fundamentally resolving the Advantage Collapsing.
-These findings highlight the need for a dynamic mechanism that can adaptively identify and select valuable rollouts to improve training efficiency.
+Our probe analysis reveals that, contrary to the ideal scenario, most rollouts exhibit advantages sharply concentrated around zero in standard RL paradigm, leading to the Advantage Collapsing phenomenon.
+This concentrated distribution weakens gradient signals, as only a few rollouts with high-magnitude advantages drive meaningful updates.
+While simply increasing the number of rollouts can partially mitigate this issue by increasing the chance of sampling valuable trajectories (Fig.
+3(a)), it substantially increases computational overhead without addressing the root cause.
+These findings highlight the need for a dynamic mechanism that adaptively selects valuable rollouts to improve training efficiency.
+Figure 3: (a) Model accuracy improves with larger rollout sizes.
+(b) Queries with different difficulties have varying accuracy during training, yielding rollouts with different diversity and qualities.
 Rollout Silencing.
-We observe that the proportion of rollouts contributing non-zero gradients steadily decreases as training progresses, resulting in the Rollout Silencing phenomenon.
-This is primarily caused by the accumulation of factors such as zero advantages, gradient clipping, and overlong truncation, which further exposes the limitations of static sampling paradigm.
-By tracking queries of varying difficulty, we find that simple queries tend to converge early, while difficult ones consistently yield low accuracy (Fig.
-2(b)).
-Both cases fail to generate diverse and informative rollouts throughout training, contributing to Rollout Silencing.
-Moreover, in standard training pipelines, each rollout is used only once for a single gradient update, limiting the model to fully exploit valuable rollouts.
-Consequently, a dynamic sampling strategy that can adaptively discard ineffective rollouts and queries, while allowing repeated learning from valuable rollouts is of great importance.
-Figure 2: (a) Model accuracy improves with larger rollout sizes.
-(b) Queries with different difficulties demonstrate varying accuracy during training, their corresponding rollouts have different diversity and qualities consequently.
-Figure 3: Overview of our proposed Shuffle-R1.
-After advantage calculation, we first conduct Pairwise Trajectory Sampling to obtain valuable trajectory pairs from original rollout pool, then perform Advantage-based Batch Shuffle to reshape the distribution of valid trajectories in a batch.
+We further observe a notable Rollout Silencing phenomenon, where the fraction of rollouts contributing non-zero gradients notably declines during training.
+It arises from factors such as zero advantages, gradient clipping, and excessive truncation, exposing the limits of static sampling paradigm.
+Tracking accuracy of queries in different difficulty reveals that simple queries converge early while difficult ones remain inaccurate (Fig.
+3(b)), both failing to generate informative rollouts.
+Moreover, standard pipelines use each rollout only once, preventing full exploitation of valuable data.
+Consequently, it is crucial to design a dynamic strategy that discards ineffective rollouts while reusing informative ones.
 3.3 Pairwise Trajectory Sampling
-To relieve Advantage Collapsing, we seek to select trajectories that offer stronger learning signals.
-Since increasing the rollout number enlarges the chance of capturing high-advantage samples, we propose Pairwise Trajectory Sampling (PTS), a data-centric module to selectively amplify valuable learning signals.
+To mitigate Advantage Collapsing, we seek to select trajectories that offer stronger learning signals.
+Inspired by the observation that a larger rollout pool increases the probability of capturing high-advantage samples, we propose Pairwise Trajectory Sampling (PTS), a data-centric module to selectively amplify valuable learning signals.
 Rather than evaluating trajectories in isolation, PTS organizes candidate rollouts into structured contrastive pairs.
@@
 The corresponding reward and advantage sets are $R=\{r_{i}\}_{i=1}^{2N}$ and $A=\{\hat{A}_{i}\}_{i=1}^{2N}$, respectively.
-To identify informative trajectory pairs, our proposed pairing mechanism follows a simple ‘max-and-min’ principle by matching the trajectory with the highest advantage to that with the lowest, the second highest to the second lowest, and so on.
+To identify informative trajectory pairs, our proposed pairing mechanism follows a straightforward ‘max-min’ pairing principle by matching the trajectory with the highest advantage to that with the lowest, the second highest to the second lowest, and so on.
 We denote the sorted advantage values in descending order as:
@@
 In this scheme, the original $2N$ rollouts are easily sorted and reorganized into $N$ pairs.
-The top-ranked pairs typically consist of trajectories with high-magnitude but opposite-sign advantages, forming contrastive pairs akin to “positive-negative” samples.
+The top-ranked pairs typically consist of trajectories with high-magnitude but opposite-sign advantages, forming contrastive pairs akin to ‘positive-negative’ samples.
 In contrast, the bottom-ranked pairs involve trajectories with advantages closer to zero.
@@
 $P_{v}=\{(o_{(i)},o_{(2N-i+1)})\}_{i=1}^{M},\quad M=\alpha N,\ \alpha\in(0,1).$ (5)
-where $\alpha$ is a hyperparameter controlling the sampling ratio from the pairing set. We only keep the trajectories in the valid set for gradient update.
+where $\alpha$ is a hyperparameter controlling the sampling ratio from the pairing set. Only the trajectories within this valid set $P_{v}$ are used for the subsequent gradient update.
 By introducing a structured contrastive sampling scheme, PTS enables more effective trajectory selection from a broader exploration space without increasing the gradient computation cost.
-The contrastive pairing not only filters out low-signal trajectories, but also sharpen the model’s policy gradient through direct comparison.
+The contrastive structure not only filters out low-signal trajectories, but also sharpen the model’s policy gradient through direct comparison.
 PTS shifts the focus of RL fine-tuning from uniform exploration to gradient-informed selection, representing a principled step toward more efficient data usage in RL training.
+Figure 4: Overview of our proposed Shuffle-R1.
+After advantage calculation, we first conduct Pairwise Trajectory Sampling to obtain valuable trajectory pairs from original rollout pool, then perform Advantage-based Batch Shuffle to reshape the distribution of valid trajectories in a batch.
 3.4 Advantage-based Batch Shuffle
-While PTS could mitigate Advantage Collapsing and improve RL training performance, the Rollout Silencing issue remains unresolved.
+While PTS mitigates Advantage Collapsing and improve RL training performance, the Rollout Silencing issue remains unresolved.
 To overcome this issue, we propose Advantage-based Batch Shuffle (ABS) module that dynamically reshapes training batches to prioritize and reinforce high-value samples.
@@
 Built on top of PTS, it serves to magnify the gradient exposure of informative samples, reshaping the training data distribution to achieve better data utilization and training efficiency.
-Denote a data batch:
-$B=\{p_{i}^{g}=(o_{i,1}^{g},\hat{A}_{i,1}^{g},o_{i,2}^{g},\hat{A}_{i,2}^{g},q^{g})\}_{i=1\sim M,g=1\sim G},$ (7)
+Denote a data batch provided by PTS:
+$B=\{p_{i}^{g}:(o_{i,1}^{g},\hat{A}_{i,1}^{g},o_{i,2}^{g},\hat{A}_{i,2}^{g},q^{g})\}_{i=1\sim M,g=1\sim G},$ (6)
 with batch size of $M\times G$, In the standard gradient update process, $B$ is sequentially divided into $K$ mini-batches, each contains $MG/K$ samples.
-In our ABS module, we first assign an importance weight to each trajectory pair based on the sum of the absolute advantages:
+In our ABS module, we first assign an importance weight to each pair $p_{j}\in B$ based on the sum of the absolute advantages:
 $W(p_{j})=|\hat{A}_{j,1}|+|\hat{A}_{j,2}|.$ (7)
-We then define the sampling probability of each pair $p_{j}\in P_{v}$ as:
-$\Phi(p_{j})=\frac{W(p_{j})}{{\sum_{k=1}^{|P_{v}|}}W(p_{k})}.$ (9)
-Based on the sampling probability, we perform $S$ sub-sampling from original batch $B$, each sub-sampling has a capacity of $T$ pairs ($2T$ trajectories):
+These weights are then normalized to form a sampling distribution $\Phi$ over the entire batch $B$:
+$\Phi(p_{j})=\frac{W(p_{j})}{{\sum_{k=1}^{|B|}}W(p_{k})}.$ (8)
+Based on the sampling distribution, we perform $S$ sub-sampling from original batch $B$, each sub-sampling has a capacity of $T$ pairs ($2T$ trajectories):
 $B_{s}=\{p_{s,t}\}_{t=1}^{T},\quad\text{s.t. }p_{s,t}\neq p_{s,t^{\prime}},\forall\,t\neq t^{\prime}.$ (9)
@@
 The reshuffled batch will maintain the gradient update paradigm of the original method.
-The ABS module optimizes the learning process through Advantage-aware shuffling and Inter-sub-batch resampling.
+The ABS module optimizes the learning process through Advantage-aware Shuffling and Sub-batch Resampling.
 It increases the update frequency of trajectories with higher advantages, maintains diversity while reinforcing high-value samples through repeated exposure.
@@
 Datasets and Benchmarks.
-We first conduct our experiments on Geometry3K dataset (Lu et al.
-2021) ($2.1k$ training samples.) and a subset of MMK12 dataset (Meng et al.
-2025) containing the same amount of data, to investigate model performance on limited training resources.
-Then, we conduct experiments on larger data scale with MM-Eureka dataset (Meng et al.
-2025), seeking for better model performance.
-We randomly select $27k$ samples from MM-Eureka dataset and mix them with Geometry3K dataset, forming a total of $30k$ training corpora.
+We first conduct our experiments on Geometry3K dataset (Lu et al., 2021) ($2.1k$ training samples, ‘Geo3K’ for short) and a subset of MMK12 dataset (Meng et al., 2025) containing the same amount of data (‘K12’ for short), to investigate model performance on limited training resources.
+To assess the scalability and effectiveness on a larger corpus, we then conduct experiments with MM-Eureka dataset (Meng et al., 2025).
+Specifically, we construct a $30k$-sample training set by combining the full Geo3K dataset with $27k$ randomly selected samples from MM-Eureka.
 All training samples are in free-form format.
 We first perform evaluation on in-domain test set of Geometry3K and MMK12.
-Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al.
-2024b), MathVision (Wang et al.
-2024), WeMath (Qiao et al.
-2024), MathVista (Lu et al.
-2023), HallusionBench (Guan et al.
-2024) and ChartQA (Masry et al.
-2022).
+Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024), MathVista (Lu et al., 2023), HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022).
 These benchmarks span across math reasoning, visual perception, and chart understanding.
-We use MathRuler to evaluate questions with free-form ground truths and Gemini-2.0-Flash-001 (Deepmind 2025) to evaluate questions with multi-choice ground truths.
+We use MathRuler to evaluate questions with free-form ground truths and Gemini-2.0-Flash-001 (Deepmind, 2025) to evaluate questions with multi-choice ground truths.
+More information in Appendix E.
+Table 1: Performance of Shuffle-R1 trained on Geometry3K dataset.
 Method Geo3K Math Avg. HallBench ChartQA
@@
 + DAPO 45.09 48.08 63.24 76.70
-+ Ours 47.88(+22.09) 48.70(+6.99) 63.19(+3.36) 77.04(+3.06)
++ GSPO 43.22 47.26 63.67 75.12
+[…]
 Qwen-7B 38.12 49.82 65.19 79.84
@@
 + DAPO 54.43 54.19 69.29 81.20
-+ Ours 55.89(+17.77) 54.63(+4.81) 69.51(+4.32) 81.64(+1.80)
-Table 1: Performance of Shuffle-R1 on Geometry3K dataset compared with GRPO and DAPO.
++ GSPO 52.83 54.27 69.48 80.96
+[…]
+Table 2: Performance of Shuffle-R1 trained on K12 dataset.
 Method K12 Math Avg. HallBench ChartQA
@@
 + DAPO 61.42 49.75 65.08 77.00
-+ Ours 62.22(+19.80) 50.05(+8.34) 65.72(+5.89) 78.28(+5.20)
++ GSPO 60.44 48.76 64.14 77.56
+[…]
 Qwen-7B 52.13 49.82 65.19 79.84
@@
 + DAPO 68.35 54.52 68.66 82.52
-+ Ours 68.78(+16.65) 55.02(+5.20) 69.87(+4.68) 82.60(+2.76)
-Table 2: Performance of Shuffle-R1 on K12 dataset compared with GRPO and DAPO.
++ GSPO 67.44 54.77 69.13 82.04
+[…]
 Implementation Details.
-We use EasyR1 (Yaowei et al.
-2025) as our training codebase.
-We choose Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct (Bai et al.
-2025) as base model to verify our method’s generalizability on model scales.
+We use EasyR1 (Yaowei et al., 2025) as our training codebase.
+We employ Qwen2.5-VL-3B/7B-Instruct (Bai et al., 2025) as base model to verify our method’s generalizability on model scales.
 Parameters of vision encoder are kept frozen.
@@
 Rollout temperature is set to 1.0 and learning rate is set to 1e61e-6.
-All experiments are conducted on $8\times$ 80G GPUs.
-For PTS, we sample 4 trajectory pairs (8 trajectories) from 16 rollouts for each query, striking a balance between training cost and exploration space.
-For ABS, we set the sub-sampling batch size ($T$) to 256 pairs (512 query-response trajectories) and the shuffle times ($S$) to 8.
-For evaluation, we set the temperature to 0.5 and report the average pass@1 accuracy of 8 tests to reduce randomness.
+All experiments are conducted on $8\times$ H800-80G GPUs.
+For each query, we generate $2N=16$ rollouts.
+For PTS, we construct $N=8$ pairs and select the top $M=4$ pairs, corresponding to a retention ratio of $\alpha=0.5$, striking a balance between training cost and exploration space.
+For ABS, we set $T=256$ pairs (512 query-response trajectories) for each sub-sampling batch.
+We construct shuffled batch by performing $S=8$ rounds of shuffle.
+Decoding temperature for evaluation is set to 0.5, and we report average pass@1 accuracy of 8 tests to reduce randomness.
 4.2 Main Results
-Model MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
-Close-source
-GPT-4o (Achiam et al. 2024) 50.8 30.4 63.8 68.8 55.0 - -
-o1 (OpenAI 2024) 57.0 60.3 73.9 - - - -
-Gemini-2.0 pro (Deepmind 2025) 67.3 48.1 71.3 - 49.8 - -
-Claude-3.7-Sonnet (Anthropic 2025) 52.0 41.3 66.8 72.6 55.4 - -
-Open-Source SFT
-InternVL-2.5-8B (Chen et al. 2025b) 39.5 17.0 64.5 - 50.1 79.1 -
-InternVL-3-8B (Zhu et al. 2025) - 29.3 71.6 - 49.9 86.6 -
-Qwen2.5-VL-3B (Bai et al. 2025) 34.8 21.9 58.4 51.7 59.8 73.1 49.9
-Qwen2.5-VL-7B (Bai et al. 2025) 42.6 25.8 67.4 63.5 65.2 79.8 57.4
-Cold Start + RL
-R1-VL-7B (Zhang et al. 2025) 40.1 24.3 62.3 59.8 60.9 76.1 53.9
-Vision-R1-7B (Huang et al. 2025) 46.1 - 70.8 - 57.8 83.1 -
-R1-OneVision-7B (Yang et al. 2025) 43.0 24.8 61.2 60.6 66.4 77.8 55.2
-OpenVLThinker-7B (Deng et al. 2025) 46.4 24.8 69.7 67.2 59.1 78.4 57.6
-VLAA-Thinker-7B (Chen et al. 2025a) 48.9 26.3 69.9 67.7 67.5 80.1 60.1
-Zero RL
-$M$-Eureka-Qwen-7B (Meng et al. 2025) 49.6 27.4 70.6 67.4 66.7 79.0 60.1
-MMR1-Math-7B (Leng et al. 2025) 39.2 31.9 71.5 70.7 69.6 82.0 60.8
-ThinkLite-VL-7B (Wang et al. 2025b) 45.2 28.0 72.4 69.3 70.2 82.0 61.2
-VL-Rethinker-7B (Wang et al. 2025a) 51.7 29.7 72.0 70.1 69.9 79.0 62.1
-NoisyRollout-7B-K12 (Liu et al. 2025a) 50.1 28.0 70.9 70.8 70.1 81.4 62.1
-Shuffle-R1-Qwen-3B (Ours) 44.2 26.8 70.4 66.5 69.2 79.9 59.5
-Shuffle-R1-Qwen-7B (Ours) 53.9 30.0 77.0 72.3 71.0 84.1 64.7
 Table 3: Model performance on representative visual reasoning benchmarks.
@@
 Best performance of RL-only models marked with Bold, second best with underline.
+Model MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
+Close-source
+GPT-4o (Achiam et al., 2024) 50.8 30.4 63.8 68.8 55.0 - -
+o1 (OpenAI, 2024) 57.0 60.3 73.9 - - - -
+Gemini-2.0 pro (Deepmind, 2025) 67.3 48.1 71.3 - 49.8 - -
+Claude-3.7-Sonnet (Anthropic, 2025) 52.0 41.3 66.8 72.6 55.4 - -
+Open-Source SFT
+InternVL-2.5-8B (Chen et al., 2025b) 39.5 17.0 64.5 - 50.1 79.1 -
+InternVL-3-8B (Zhu et al., 2025) - 29.3 71.6 - 49.9 86.6 -
+Qwen2.5-VL-3B (Bai et al., 2025) 34.8 21.9 58.4 51.7 59.8 73.1 49.9
+Qwen2.5-VL-7B (Bai et al., 2025) 42.6 25.8 67.4 63.5 65.2 79.8 57.4
+Cold-Start + RL
+R1-VL-7B (Zhang et al., 2025) 40.1 24.3 62.3 59.8 60.9 76.1 53.9
+Vision-R1-7B (Huang et al., 2025) 46.1 - 70.8 - 57.8 83.1 -
+R1-OneVision-7B (Yang et al., 2025) 43.0 24.8 61.2 60.6 66.4 77.8 55.2
+OpenVLThinker-7B (Deng et al., 2025) 46.4 24.8 69.7 67.2 59.1 78.4 57.6
+VLAA-Thinker-7B (Chen et al., 2025a) 48.9 26.3 69.9 67.7 67.5 80.1 60.1
+Zero RL
+$M$-Eureka-Qwen-7B (Meng et al., 2025) 49.6 27.4 70.6 67.4 66.7 79.0 60.1
+MMR1-Math-7B (Leng et al., 2025) 39.2 31.9 71.5 70.7 69.6 82.0 60.8
+ThinkLite-VL-7B (Wang et al., 2025b) 45.2 28.0 72.4 69.3 70.2 82.0 61.2
+VL-Rethinker-7B (Wang et al., 2025a) 51.7 29.7 72.0 70.1 69.9 79.0 62.1
+NoisyRollout-7B-K12 (Liu et al., 2025a) 50.1 28.0 70.9 70.8 70.1 81.4 62.1
+[…]
+[…]
 Comparison with Representative Algorithms.
-We compare model performance of our method with GRPO (Shao et al.
-2024) and DAPO (Yu et al.
-2025).
-As shown in Tab.
-1, after training on the Geometry3K dataset, the 3B model trained with our method achieved 47.88% accuracy on in-domain test set, with a 5.2% improvement compared to model trained with GRPO and 2.7% improvement compared to DAPO.
-Similarly, for the 7B model, our method led to 3.3% improvement over GRPO and 1.4% over DAPO, achieving 55.89% accuracy.
-The superiority of our framework is further highlighted in out-of-domain benchmarks.
-For out-of-domain math reasoning task, the average accuracy is improved by 1.96% on 3B model and 1.5% on 7B model compared to GRPO, and surpasses DAPO as well.
-Similar performance gains are also observed on HallusionBench and ChartQA, where the data gap is even further from training data, showing strong generalizability of our framework.
-Tab.
-2 demonstrates similar performance results on K12 dataset, where consistent performance gains are achieved on both in-domain and out-of-domain benchmarks, highlighting our framework’s generalizability on different data distribution.
+We compare our method with GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025).
+On Geometry3K (Tab.
+2), our 3B model reaches 47.88% accuracy, outperforming GRPO by 5.2% and DAPO by 2.7%.; the 7B model achieves 55.89%, with gains of 3.3% and 1.4%, respectively.
+On out-of-domain math reasoning tasks, our method further improves average accuracy by 1.96% (3B) and 1.5% (7B) over GRPO, also surpassing DAPO.
+Consistent gains are observed on HallusionBench and ChartQA, where data distribution diverges more from training.
+Similar trends hold on the K12 experiments (Tab.
+2), with improvements in both in-domain and out-of-domain settings, underscoring the robustness of our framework across tasks and distributions.
+Compared with the latest GSPO (Zheng et al., 2025), which replaces token-level importance sampling with sequence-level importance sampling, our method also demonstrates superior performance on both in-domain and out-of-domain tasks.
+All the performance results above indicate our framework’s generalizability on different data distributions and model scales, highlighting the effectiveness of dynamic RL training paradigm.
+More results in Appendix C.
+Figure 5: Advantage distribution in a training batch of GRPO and our framework.
 Comparison with RL-based models.
@@
 MM-Eureka with direct RL and VLAA-Thinker with RL after cold-start SFT.
-Notably, our model achieves competitive or superior performance on several benchmarks compared to leading close-source models, for instance, Claude-3.7-Sonnet (Anthropic 2025) and GPT-4o (Achiam et al.
-2024).
+Notably, our model achieves competitive or superior performance on several benchmarks compared to leading close-source models, for instance, Claude-3.7-Sonnet (Anthropic, 2025) and GPT-4o (Achiam et al., 2024).
 Under the same setting, our 3B variant also demonstrates strong performance, even outperforming several 7B models on certain benchmarks.
 These results highlight the superiority of our proposed approach in boosting the training efficiency in reinforcement learning.
-Figure 4: Advantage distribution in a training batch of GRPO and our framework.
 Efficiency Analysis.
 The improvement in model performance mainly stems from better training efficiency.
-Fig.
-2(a) in previous section shows that, under the same update rollout size, model trained under our framework significantly outperforms GRPO, indicating more effective utilization of training data.
 Advantage distribution analysis in Fig.
-4 proves that, PTS effectively mitigates Advantage Collapsing by increasing the proportion of large-magnitude advantages.
+5 confirms that, PTS effectively mitigates Advantage Collapsing by increasing the proportion of large-magnitude advantages.
 ABS further optimizes the batch composition, enabling the model to focus on more informative trajectories.
 Fig.
-5(a) and (b) further probe into training dynamics and demonstrate that our framework consistently achieves higher training and validation accuracy, reaching comparable performance as GRPO but with only half the training steps.
+7(a) and (b) further probe into training dynamics and demonstrate that our framework consistently achieves higher training and validation accuracy, reaching comparable performance as GRPO with as little as half the training steps.
 Moreover, our framework effectively mitigates the issue of “Rollout Silencing” shown in Fig.
-5(c), maintaining a high token utilization rate across all training stages.
+7(c), maintaining a high token utilization rate across all training stages.
 Fig.
-5(d) further illustrates the favorable trade-off between training scale and computational cost of our approach, substantially expanding the RL exploration space with minimal additional overhead.
+7(d) further illustrates the favorable trade-off between training scale and computational cost of our approach, uplifting training performance by a large margin.
+Fig.
+7 illustrates the actual wall-clock GPU time between GRPO and Shuffle-R1.
+Under the same update size and total training steps, Shuffle-R1 achieves substantially higher train/val accuracy than GRPO in the early training stage.
+More importantly, the total GPU time of Shuffle-R1 only increases by 4% $\sim$ 7.7% relative to GRPO.
+When targeting the same accuracy as GRPO, Shuffle-R1 requires roughly half the number of training steps and approximately 60% of the total wall-clock time.
 4.3 Ablation Study
-Figure 5: (a): Training accuracy of GRPO and Shuffle-R1.
-(b): Validation accuracy of GRPO and Shuffle-R1.
-(c): Token utilization rate of GRPO and Shuffle-R1.
-(d): Shuffle-R1 achieves better performance with minimal extra time cost.
+Figure 6: (a) Training accuracy of Shuffle-R1 on Geo3K.
+(b) Validation accuracy of Shuffle-R1 on Geo3K.
+(c) Token utilization rate of Shuffle-R1 on Geo3K.
+(d) Shuffle-R1 achieves better performance with minimal extra time cost.
+Figure 7: Wall-Clock Training Curve of Shuffle-R1 compared with GRPO.
 We conduct ablation experiments on Qwen2.5-VL-3B-Instruct using Geometry3K, focusing on two objectives: (1) assessing the contribution of each component in our framework, and (2) validating the impact of key designs.
+More detailed results are reported in Appendix C.
+Component-wise Contribution.
+We evaluate the effectiveness of PTS and ABS by incrementally adding them to the baseline.
+As shown in Tab.
+5, On in-domain Geometry3K test set, PTS lifts accuracy from 42.64% to 46.21% (+3.57).
+Adding ABS yields a further +1.67, reaching 47.88%.
+Similar improvement trends appear on out-of-domain benchmarks: on the math reasoning set, the full setting (PTS + ABS) attains 48.70% vs.
+46.74% with GRPO (41.71% before RL training); on ChartQA, we also observe effective performance gain (77.04% vs.
+76.20%, with 73.08% before RL training) despite larger distribution shift.
+Our method also demonstrates improved performance in HallusionBench.
+To isolate the contribution of ABS, we conducted an experiment where ABS is applied to trajectory pairs formed by uniform random sampling instead of PTS.
+This setting (Tab.
+5 line 4) also demonstrates notable improvement over the baseline, confirming its effectiveness.
+Table 4: Ablation on effectiveness of PTS and ABS.
 GRPO PTS ABS Geo3k Math Avg. HallBench ChartQA
@@
 46.21 47.64 63.40 76.52
-47.88 48.70 63.19 77.04
-Table 4: Ablation study on effectiveness of PTS and ABS.
-Effectiveness of Algorithm Design.
-We assess the effectiveness of proposed PTS and ABS modules by successively introduce them to the baseline.
-As shown in Tab.
-4, in the in-domain Geometry3K test set, PTS demonstrates significant performance gains, improving the model accuracy from 42.64% to 46.21% (+3.57%).
-When ABS is introduced for co-optimization, the model achieves a further 1.67% performance gain, eventually reaching an accuracy of 47.88%.
-This improvement trend continues in the out-of-domain evaluation scenario: in the math reasoning task, the full setting (PTS + ABS) achieves an absolute performance improvement of 2.0% compared to GRPO (48.70% vs.
-46.74%, with 41.71% before post-training).
-In the ChartQA dataset, which has more significant distributional differences, the joint application of PTS and ABS still maintains an effective performance gain (77.04% vs.
-76.20%, with 73.08% before post-training).
-Our method also demonstrates improved performance in HallusionBench.
-Rationality of Pairwise Trajectory Sampling.
-One core mechanism of the PTS lies in the structured contrastive sampling scheme.
-To verify the effectiveness of this bidirectional optimization mechanism, we designed the following experiments: (1) one-way positive sampling (only select trajectories with highest advantage); (2) one-way negative sampling (only select trajectories with lowest advantage); and (3) unbiased random sampling.
-All settings maintained a consistent sampling ratio (8 valid trajectories from 16 rollouts) to ensure fairness.
-We disable ABS since it relies on the pairing result of PTS.
-As shown in Tab.
-5, model trained with PTS receives a consistent performance gain on both in-domain and out-of-domain tasks, while one-way positive and one-way negative sampling show a decline in all tasks, demonstrating the effectiveness and rationality of our design.
-Unbiased random sampling only receives minor improvement over baseline, far behind the effectiveness of PTS.
+46.82 47.79 62.77 75.12
+[…]
+Table 5: Ablation on rationality of PTS and ABS.
 Setting Geo3k Math Avg. HallBench ChartQA
@@
 + only min 23.36 41.52 60.98 74.36
-+ rand. select 43.53 46.62 63.19 76.00
-+ PTS 46.21 47.64 63.40 76.52
++ random pick 43.53 46.62 63.19 76.00
+[…]
 Ablation on ABS
-+ rand. shuffle 46.05 47.40 63.19 76.60
++ random shuffle 46.05 47.40 63.19 76.60
 + reorder 46.28 47.64 63.09 76.64
-+ ABS 47.88 48.80 63.19 77.04
-Table 5: Ablation study on rationality of PTS and ABS.
-Rationality of Advantage-based Batch Shuffle.
+[…]
+Analysis of Pairwise Trajectory Sampling.
+One core mechanism of the PTS lies in the structured contrastive sampling scheme.
+To validate the effectiveness of our bidirectional, contrastive sampling scheme, we compare PTS against three alternative strategies: (1) One-way Positive Sampling (only select trajectories with highest advantage); (2) One-way Negative Sampling (only select trajectories with lowest advantage); and (3) Unbiased Random Sampling.
+All settings maintained a consistent sampling ratio (8 valid trajectories from 16 rollouts) to ensure fairness.
+We disable ABS since it relies on the pairing result of PTS.
+As shown in Tab.
+5, model trained with PTS receives a consistent performance gain on both in-domain and out-of-domain tasks, while both one-way positive and negative sampling result in a performance decline even below the GRPO baseline, demonstrating the effectiveness and rationality of our design.
+Unbiased random sampling only receives minor improvement over baseline, far behind the effectiveness of PTS.
+Figure 8: Ablation on key hyper parameters. (a) Effect of different sampling ratio $\alpha$. (b) Effect of different shuffle times $S$.
+Analysis of Advantage-based Batch Shuffle.
 ABS introduces Advantage-aware shuffling and Inter-sub-batch resampling to reshape the training batch.
 To validate their effectiveness, we designed two contrastive experiments: (1) Unbiased Shuffle: using uniformly distributed sampling weights to perform shuffle strategy.
-and (2) static reorder: randomly reorder the training batch without shuffle, maintaining the original data distribution.
+and (2) Static Reorder: randomly reorder the training batch without sub-sampling, maintaining the original data distribution.
 We enable PTS during training as ABS relies on its pairing result.
@@
 5, model trained with ABS significantly outperforms the contrastive settings.
-The unbiased shuffle setting even performs worse compared to the PTS-only setting, demonstrating the rationality of ABS.
+The unbiased shuffle setting even performs worse compared to the PTS-only setting, demonstrating significance of advantage-weighting.
 The static reorder setting has no improvement compared to PTS-only setting, as they have the same data distribution.
@@
 As shown in Fig.
-6(a), both $\alpha=0.75$ and $\alpha=0.5$ yield strong performance, while $\alpha=0.25$ lags behind.
+8(a), both $\alpha=0.75$ and $\alpha=0.5$ yield strong performance, while $\alpha=0.25$ lags behind.
 We attribute this to over-pruning, where filtering out rollouts too aggressively may reduce data diversity.
-We choose $\alpha=0.5$ to strike a balance between signal quality and computational efficiency.
+We choose $\alpha=0.5$ for a balance between signal quality and computational efficiency.
 For $S$, we fix $\alpha=0.5$ and vary the shuffle times as 4, 8, and 16.
 Fig.
-6(b) shows that performance improves with increasing $S$, but saturates beyond $S=8$.
+8(b) shows that performance improves with increasing $S$, but saturates beyond $S=8$.
 This suggests that moderate resampling enhances data exposure, but too many shuffles may offer diminishing returns.
-Figure 6: Ablation on key hyper parameters. (a): Effect of sampling ratio $\alpha$. (b): Effect of shuffle times $S$.
+More details in Appendix C.
+4.4 Extension Experimental Results
+Table 6: Extension experiments on Qwen2.5-VL-32B.
+Model MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
+Qwen2.5-32B 57.0 38.2 75.4 72.9 71.3 80.7 65.9
++ GRPO 58.4 39.3 77.3 75.9 70.3 83.8 67.4
+[…]
+Table 7: Extension experiments on Referring Expression Comprehension task.
+Model RefCOCO(testA) RefCOCO(testB) RefCOCO+(testA) RefCOCO+(testB) RefCOCOg(test)
+Qwen2.5-3B 86.09 75.64 81.71 66.93 72.39
++GRPO 89.90 81.33 85.94 70.97 81.45
+[…]
+Scaling to 32B Model
+We conduct experiments on Qwen2.5-VL-32B to inveritage the scalability and generalizability of Shuffle-R1 on larger models.
+We train the 32B variant on the selected $30k$ data for 50 steps (to save training resources cost).
+Results in Tab.
+7 demonstrate that Shuffle-R1 also performs well on 32B scale.
+Referring Expression Comprehension
+For tasks beyond math/visual reasoning, we conduct an experiment on Referring Expression Comprehension (REC).
+We train Qwen2.5-VL-3B on 60K data randomly selected from RefCOCO/RefCOCOg/RefCOCO+ using a IoU-based soft reward.
+Results in Tab.
+7 show that Shuffle-R1 can be well adapted to soft-reward tasks and beyond.
+Table 8: Extension experiments on LLM.
+Setting Math12K AIME24 MATH500 GSM8K GPQA Olymp.
+Qwen2.5-Math-1.5B-Base
++ GRPO 67.6 10.0 66.0 74.0 25.7 33.3
+[…]
+Qwen2.5-Math-7B-Base
++ GRPO 74.6 20.0 76.4 84.8 36.3 39.7
+[…]
+Language-only Reasoning
+While our initial analysis focused on the RL training dynamics of MLLMs, we conduct an early extension experiment to validate the feasibility of Shuffle-R1 on LLMs.
+We train Qwen2.5-Math-1.5B/7B-Base (Yang et al., 2024) on Open-S1 dataset (Dang and Ngo, 2025) for 150 steps and evaluate model performance on Math12K (Hendrycks et al., 2021), AIME24, MATH500 (Lightman et al., 2023), GSM8K (Cobbe et al., 2021), GPQA-Diamond (Rein et al., 2023) and OlympiadBench (He et al., 2024).
+As shown in Tab.
+8, our framework delivers significant improvements on all the benchmarks compared to GRPO, demonstrating its potential effectiveness on text-only LLMs.
 5 Conclusion
-In this paper, we propose Shuffle-R1, a simple but effective framework that improves the training efficiency of reinforcement learning of multimodal large language models.
+In this paper, we introduce Shuffle-R1, a simple but effective framework designed to improve the training efficiency of reinforcement learning of multimodal large language models.
 Through Pairwise Trajectory Sampling and Advantage-based Batch Shuffle, our framework significantly outperforms representative algorithms and models in both in-domain and out-of-domain tasks, demonstrating the value of data-centric adaptive design.
 We hope that our motivations, method, and findings are helpful for further research.
+Acknowledgment
+This work was done during the research internship of Linghao Zhu, Yiran Guan, and Dingkang Liang at Xiaomi.
+This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62225603 and 623B2038.
+Ethics statement
+All training and evaluation data used in this project are collected from public academic datasets.
+All the pre-trained model checkpoints and code framework employed are forked from open-source research projects.
+During the research process, we strictly adhered to the open-source licenses and usage requirements of all relevant parties.
+No non-public / private data was used.
+Furthermore, all authors strictly complied with the ICLR Code of Ethics throughout the research process and affirm that no activities violating these ethical standards were involved.
+Reproducibility statement
+We have reported the base model checkpoints, code framework, training data, evaluation benchmarks and metrics, and core hyperparameter settings used in the project in the main paper.
+A detailed pseudo code explaining our framework flow is provided in Appendix A.
+More detailed hyperparameter settings and information about the training infrastructure are provided in Appendix D.
+In addition, we have also included the prompts used during training and inference, examples of the training and evaluation data, as well as examples of the model inference results in Appendix EF and G.
+LLM Usage Statement
+During this research process, the usage of LLM is strictly restricted and was used in limited ways: (1) GitHub Copilot was employed during code development for code checking and bug fixes; (2) LLM-based tools were used during paper writing for language refinement and grammar correction, and (3) LLM judges were applied as a part of evaluation metrics, including key answer extraction and correctness assessment.
+All core components of this work, including research motivation, algorithm design, core code implementation, data collection, and experiment analysis, and experimental conclusions, were conducted entirely and independently by the authors without reliance on LLMs.
 Appendix A Pseudo Code
@@
 […]
-Appendix C Experiment Settings
+Appendix C More Experimental Results
+Detailed Model Performance.
+We provide a more detailed performance of models trained on Geometry3K and K12 dataset in Tab.
+9, reporting model performance on each out-of-domain benchmarks as a supplement to “Math Avg.” columns in main paper.
+Trained with only $2.1k$ data, both the 3B and 7B model demonstrate significant performance gains.
+Full Setting Comparison.
+Improved training algorithm leads to improved performance.
+To better reveal the model performance beyond the limited data scale of Geo3K, we conduct full 30k-sample 150-step experiment on Qwen2.5-VL-7B.
+As shown in Tab.
+10, Shuffle-R1 achieves clear and consistent gains over strong baselines across all six evaluation benchmarks.
+This result further proven the robustness and scalability of Shuffle-R1.
+Comparison with more RL algorithms.
+We provide additional comparisons with more representative RL algorithms, i.e.
+RLOO (Ahmadian et al., 2024) and REINFORCE++ (Hu, 2025).
+We conduct experiments on Qwen2.5-VL-3B-Instruct with Geometry3K dataset.
+All the experiment settings are kept the same as GRPO/DAPO/GSPO in main paper.
+The final model performance is shown in Tab.
+11.
+Our framework outperform these algorithms by a large margin in both in-domain and out-of-domain tasks.
+Figure 9: Performance of Shuffle-R1 compared with directly applying Prioritized Experience Replay.
+Table 9: Detailed performance on out-of-domain benchmarks of models trained on Geometry3K and K12 data. Highest accuracy marked in Bold.
+Method MathVerse MathVision MathVista WeMath HallBench ChartQA Total Avg.
+Qwen2.5-VL-3B 34.77 21.94 58.40 51.72 59.83 73.08 49.96
+[…]
+[…]
+Qwen2.5-VL-7B 42.59 25.76 67.40 63.51 65.19 79.84 57.38
+[…]
+[…]
+Table 10: Full 30k experiment on Qwen2.5-VL-7B. Highest accuracy marked in Bold.
+Method MathVerse MathVision MathVista WeMath HallBench ChartQA Total Avg.
+Qwen2.5-VL-7B 42.6 25.8 67.4 63.5 65.2 79.8 57.4
++ GRPO 50.6 28.3 74.5 69.7 70.7 81.4 62.5
++ DAPO 51.4 28.8 75.3 71.5 71.0 82.8 63.4
++ GSPO 50.8 28.2 75.3 70.1 69.7 82.9 62.8
+[…]
+Table 11: Performance of Shuffle-R1 on Geometry3K dataset compared with RLOO and REINFORCE++. Highest accuracy marked in Bold.
+Method Geo3K MathVerse MathVision MathVista WeMath HallBench ChartQA
+Qwen2.5-VL-3B 25.79 34.77 21.94 58.40 51.72 59.83 73.08
++ RLOO 42.09 39.94 22.96 58.90 59.48 64.24 76.68
++ REINFORCE++ 41.76 41.70 24.86 60.90 63.51 62.99 76.20
+[…]
+Comparison with Prioritize Experience Replay.
+We conduct a comparative study by replacing ABS with a prioritized experience replay mechanism.
+Experience replay maintains a decoupled buffer of past samples, whereas ABS adopts an online, in-place shuffle strategy to dynamically reshape the data distribution.
+As shown in Fig.
+9(a) and (b), in the later stages, the experience replay setting exhibits a plateau in training accuracy and even a drop in validation accuracy, indicating potential overfitting to stale samples.
+This suggests that prioritized experience replay may overly emphasize historical trajectories, leading to suboptimal convergence.
+Moreover, ABS proves to be more effective in mitigating the Rollout Silencing.
+Detailed Impact of Hyperparameters.
+Here, we provide detailed model performance comparison under different hyperparameter settings in Fig.
+12, as supplementary information to corresponding ablation experiments in main paper.
+For sampling ratio ($\alpha$), both the in-domain and out-of-domain accuracy improves as we increase $\alpha$ from 0.25 to 0.5.
+The model performance shows a mild decrease when $\alpha$ is further raised to 0.75.
+This result demonstrates again that not all the rollouts contribute equally positive to RL training.
+For shuffle times ($S$), the model performance receives a consistent gain when $S$ is increased from 4 to 8, but declines when $S$ is set to 16.
+This result suggests that appropriate repeated training on high-quality samples can enhance the model’s reasoning ability without affecting data diversity.
+Table 12: Performance of Shuffle-R1 under different hyperparameter settings. Highest accuracy marked in Bold.
+Setting Geo3K MathVerse MathVision MathVista WeMath HallBench ChartQA
+Ablation on sampling ratio $\alpha$ ($S=8$)
+$\alpha=0.25$ 45.75 42.43 23.05 60.20 63.62 63.09 75.36
+[…]
+$\alpha=0.75$ 47.41 41.70 25.62 60.90 64.54 62.99 73.36
+Ablation on shuffle times $S$ ($\alpha=0.5$)
+$S=4$ 45.92 42.21 24.78 62.30 64.08 62.25 75.60
+[…]
+$S=16$ 47.23 43.70 24.86 60.90 62.64 62.20 74.84
+Appendix D Experiment Settings
 We report details of our training and evaluation settings here, including reward function design, main hyperparameters and computing resources.
-Reward Calculation.
-We adopt a combination of format reward and accuracy reward as the final reward in reinforcement learning. The format reward and accuracy reward are calculated as follows:
-$r_{\text{format}}=\begin{cases}1,&\text{if format is {correct}}\\ 0,&\text{if format is {incorrect}}\end{cases}$ (11)
-[…]
-The final reward is the weighted sum of above rewards:
-$r_{\text{overall}}=0.1\times r_{\text{format}}+0.9\times r_{\text{acc}}$ (13)
-Format reward is assigned to a smaller weight since response formatting is easy to learn.
-Hyperparameters.
-We use EasyR1 (Yaowei et al.
-2025) as our training framework.
-Full hyperparameter settings during training is shown in Tab.
-6.
-For experiments on Geometry3K and K12 with $\sim 2.1k$ training samples, we set the training steps to 80.
-For the joint training experiments ($\sim 30k$ training samples), we increase the training steps to 150 due to extended data size.
-Other hyperparameters that are not mentioned are kept to default values of EasyR1.
+Table 13: Hyperparameter settings.
 Hyperparameters Value
@@
 vision encoder frozen
-Table 6: Hyperparameter settings.
+Reward Calculation.
+We adopt a combination of format reward and accuracy reward as the final reward in reinforcement learning. The format reward and accuracy reward are calculated as follows:
+$r_{\text{format}}=\begin{cases}1,&\text{if format is {correct}}\\ 0,&\text{if format is {incorrect}}\end{cases}$ (10)
+[…]
+The final reward is the weighted sum of above rewards:
+$r_{\text{overall}}=0.1\times r_{\text{format}}+0.9\times r_{\text{acc}}$ (12)
+Format reward is assigned to a smaller weight since response formatting is easy to learn.
+Hyperparameters.
+We use EasyR1 (Yaowei et al., 2025) as our training framework.
+Full hyperparameter settings during training is shown in Tab.
+13.
+For experiments on Geometry3K and K12 with $\sim 2.1k$ training samples, we set the training steps to 80.
+For the joint training experiments ($\sim 30k$ training samples), we increase the training steps to 150 due to extended data size.
+Other hyperparameters that are not mentioned are kept to default values of EasyR1.
 Computing Resources.
-All experiments are conducted on $8\times$ 80G-SXM GPUs.
-Appendix D Dataset Collection
-For training dataset, we choose Geometry3K (Lu et al.
-2021) and MMK12 dataset.
+All experiments are conducted on $8\times$ NVIDIA H800-80G-SXM GPUs.
+Appendix E Dataset Collection
+Figure 10: Top: Examples of Geometry3K. Bottom: Examples of MMK12.
+Figure 11: Top: Examples of math reasoning tasks. Middle: Examples of visual perception tasks. Bottom: Examples of chart understanding tasks.
+For training dataset, we choose Geometry3K (Lu et al., 2021) and MMK12 dataset.
 Geometry3K dataset is a high quality real-world geometry problem solving dataset, containing 2.1K training samples and 601 test samples.
@@
 The text problems are very compact with most of the geometric conditions represented in images, making it suitable for our training.
-MMK12 dataset is introduced by MM-Eureka (Meng et al.
-2025), which contains 16k math reasoning training samples.
+MMK12 dataset is introduced by MM-Eureka (Meng et al., 2025), which contains 16k math reasoning training samples.
 The training samples have both geometric and non-geometric problems and have been carefully examined and manually filtered to ensure quality.
 The test set of MMK12 further includes other STEM problems such as physics, biology and chemistry.
-During our training, we used a randomly selected subset of MMK12 provided by NoisyRollout (Liu et al.
-2025a) (referred to as ‘K12’ in the paper to distinguish it from full set of MMK12), which has 2.1k samples with the same size as Geometry3K.
+During our training, we used a randomly selected subset of MMK12 provided by NoisyRollout (Liu et al., 2025a) (referred to as ‘K12’ in the paper to distinguish it from full set of MMK12), which has 2.1k samples with the same size as Geometry3K.
 All the training samples are in free-form format.
 Examples of training data shown in Fig.
-7.
-Figure 7: Top: Examples of Geometry3K. Bottom: Examples of MMK12.
+11.
 To fully evaluate model’s reasoning ability, apart from in-domain test set from training dataset, we select several representative benchmarks to examine model’s performance on math reasoning, visual perception, and chart understanding tasks.
-For math reasoning task, we choose MathVerse (Zhang et al.
-2024b), MathVision (Wang et al.
-2024), WeMath (Qiao et al.
-2024) and MathVista (Lu et al.
-2023).
-We choose HallusionBench (Guan et al.
-2024) and ChartQA (Masry et al.
-2022) for visual perception and chart understanding tasks respectively.
+For math reasoning task, we choose MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024) and MathVista (Lu et al., 2023).
+We choose HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022) for visual perception and chart understanding tasks respectively.
 We believe that a good RL algorithm can not only improve in-domain performance, but also have the potential to promote robust generalization to out-of-domain tasks.
 Examples of evaluation data shown in Fig.
-8.
-Figure 8: Top: Examples of math reasoning tasks. Middle: Examples of visual perception tasks. Bottom: Examples of chart understanding tasks.
-Appendix E Evaluation Metrics
+11.
+Appendix F Evaluation Metrics
 We adopt pass@1 accuracy as evaluation metric.
@@
 Evaluation settings of our model and all the reproduced results in main paper are kept the same to ensure a fair comparison.
-For evaluation of MathVerse, MathVision, WeMath, MathVista and ChartQA, we employ Gemini-2.0-Flash-001 (Deepmind 2025) to first extract the predicted answer from model response then compare it with ground truth.
+For evaluation of MathVerse, MathVision, WeMath, MathVista and ChartQA, we employ Gemini-2.0-Flash-001 (Deepmind, 2025) to first extract the predicted answer from model response then compare it with ground truth.
 Fig.
@@
 Figure 13: Score prompt for Gemini.
-Appendix F More Experimental Results
-Detailed Model Performance.
-We provide a more detailed performance of models trained on Geometry3K and K12 dataset in Tab.
-7, reporting model performance on each out-of-domain benchmarks as a supplement to “Math Avg.” columns in main paper.
-Trained with only $2.1k$ data, both the 3B and 7B model demonstrate significant performance gains.
-Comparison with more RL algorithms.
-We provide additional comparisons with more representative RL algorithms, i.e.
-RLOO (Ahmadian et al.
-2024) and REINFORCE++ (Hu 2025).
-We conduct experiments on Qwen2.5-VL-3B-Instruct with Geometry3K dataset.
-All the experiment settings are kept the same as GRPO and DAPO in main paper.
-The final model performance is shown in Tab.
-8.
-Our framework outperform these algorithms by a large margin in both in-domain and out-of-domain tasks.
-Figure 11: Performance of Shuffle-R1 compared with directly applying Prioritized Experience Replay.
-Method MathVerse MathVision MathVista WeMath HallBench ChartQA Total Avg.
-Qwen2.5-VL-3B 34.77 21.94 58.40 51.72 59.83 73.08 49.96
-+ Ours (Geo3K) 43.55 25.30 61.80 64.14 63.19 77.04 55.84
-+ Ours (K12) 44.06 26.48 64.90 64.77 65.72 78.28 57.37
-Qwen2.5-VL-7B 42.59 25.76 67.40 63.51 65.19 79.84 57.38
-+ Ours (Geo3K) 50.96 27.47 70.90 69.19 69.51 81.64 61.61
-+ Ours (K12) 48.59 28.61 73.20 69.71 69.87 82.60 62.09
-Table 7: Detailed performance on out-of-domain benchmarks of models trained on Geometry3K and K12 data. Highest accuracy marked in Bold.
-Method Geo3K MathVerse MathVision MathVista WeMath HallBench ChartQA
-Qwen2.5-VL-3B 25.79 34.77 21.94 58.40 51.72 59.83 73.08
-+ RLOO 42.09 39.94 22.96 58.90 59.48 64.24 76.68
-+ REINFORCE++ 41.76 41.70 24.86 60.90 63.51 62.99 76.20
-+ Ours 47.88 43.55 25.30 61.80 64.14 63.19 77.04
-Table 8: Performance of Shuffle-R1 on Geometry3K dataset compared with RLOO and REINFORCE++. Highest accuracy marked in Bold.
-Comparison with Direct Experience Replay.
-We conduct a comparative study by replacing ABS with a prioritized experience replay mechanism.
-Experience replay maintains a decoupled buffer of past samples, whereas ABS adopts an online, in-place shuffle strategy to dynamically reshape the data distribution.
-As shown in Fig.
-11(a) and (b), in the later stages, the experience replay setting exhibits a plateau in training accuracy and even a drop in validation accuracy, indicating potential overfitting to stale samples.
-This suggests that prioritized experience replay may overly emphasize historical trajectories, leading to suboptimal convergence.
-Moreover, ABS proves to be more effective in mitigating the Rollout Silencing.
 Appendix G Case Study
@@
 This modification has resulted in a perception error and in base model, but the RL model can accurately identify the inserted image.
+Appendix H Theoretical Analysis of Shuffle-R1
+In this section, we provide an intuitive analysis on how the adaptive selection & resampling improve the training dynamics, and discuss the bias introduced by this process.
+We begin by assuming the Advantage($A$) follows a Gaussian distribution, $A\sim\mathcal{N}(0,\sigma^{2})$, which matches the observation of our probe analysis.
+We define the original expectation of $|A|$ as:
+$E_{\text{orig}}=\mathbb{E}[|A|].$ (13)
+We simplify the PTS as a truncate process with a threshold $\tau>0$, and define the kept set $S^{\text{PTS}}=\{|A|>\tau\}$, the corresponding conditional expectation is:
+$E_{\text{PTS}}=\mathbb{E}[|A|~|~|A|\geq\tau].$ (14)
+On the truncated set, we sample with weights $w(a)\propto|a|$, and the resulting expectation is denoted as:
+$E_{\text{new}}=\mathbb{E}[w(a)|a|~|~|a|\in S^{\text{PTS}}].$ (15)
+H.1 Intuitive Analysis on Training Dynamics
+Proposition 1 (Amplification on expectation of Advantage).
+Proof. Define $X=|A|$ on the $S^{\text{PTS}}$, the weighted sampling probability density can be expressed as:
+$p_{\text{new}}(x)=\frac{x\cdot p(x~|~S^{\text{PTS}})}{\mathbb{E}[X~|~S^{\text{PTS}}]}=\frac{x\cdot p(x~|~S^{\text{PTS}})}{E_{\text{PTS}}}.$ (16)
+Thus the weighted sampling expectation can be calculated as:
+$E_{\text{new}}=\int x\cdot\frac{x\cdot p(x~|~S^{\text{PTS}})}{E_{\text{PTS}}}dx=\frac{\mathbb{E}[X^{2}|S^{\text{PTS}}]}{\mathbb{E}[X|S^{\text{PTS}}]}=\frac{E_{\text{PTS}}^{(2)}}{E_{\text{PTS}}}.$ (17)
+Based on Cauchy–Schwarz inequality, we can derive that:
+$\mathbb{E}[X^{2}]\geq(\mathbb{E}[X])^{2}\Longrightarrow E_{\text{new}}\geq E_{\text{PTS}}$ (18)
+Further, the original expectation can be decomposed as:
+$E_{\text{orig}}=(1-p_{\tau})\mathbb{E}[|A|~|~|A|<\tau]+p_{\tau}\mathbb{E}[|A|~|~|A|\geq\tau],\quad\text{where}\ p_{\tau}=\mathbb{P}(|A|\geq\tau).$ (19)
+Since
+$\mathbb{E}[|A|~|~|A|<\tau]<\tau<\mathbb{E}[|A|~|~|A|\geq\tau].$ (20)
+we have $E_{\text{PTS}}>E_{\text{orig}}\Longrightarrow E_{\text{new}}>E_{\text{orig}}$. $\qquad\square$
+Proposition 2 (Amplification on gradient norm).
+Proof. In the policy gradient training, the per-sample gradient can be expressed as:
+$g=A\cdot u.\qquad u=\nabla_{\theta}\text{log}\pi_{\theta}(o).$ (21)
+g=|A|u.||g||=|A|~||u||. (22)
+Derived from the standard assumptions of policy gradient that $u||$ is approximately independent from $|A|$, we have:
+[…]
+which in turn yields larger update magnitudes per optimization step. $\qquad\square$
+H.2 Bias Under Selective Sampling
+Proposition 3 (Adaptive sampling introduce positive bias).
+Proof. Under the Gaussian distribution assumption, the expectation can be calculated as:
+$\mathbb{E}_{\text{new}}[A]$ $=\frac{1}{Z}(\int_{-\infty}^{-\tau}a|a|p(a)da+\int_{\tau}^{\infty}a|a|p(a)da)$
+$=\frac{2}{Z}\int_{\tau}^{\infty}a^{2}p(a)da>0$
+where the normalized factor $Z$ is defined as:
+$Z$ $=\int_{-\infty}^{\infty}|a|p(a)\mathbb{I}(|a|>\tau)da$
+$=2\int_{\tau}^{\infty}a\cdot p(a)da$
+The positive bias indicates that the gradient update is systematically tend to trajectories that has higher advantage. $\qquad\square$
 Figure 14: A case study on math related task.

### v5

@@
 We first perform evaluation on in-domain test set of Geometry3K and MMK12.
-Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024), MathVista (Lu et al., 2023), HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022).
+Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2025), MathVista (Lu et al., 2023), HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022).
 These benchmarks span across math reasoning, visual perception, and chart understanding.
@@
 Rollout temperature is set to 1.0 and learning rate is set to 1e61e-6.
-All experiments are conducted on $8\times$ H800-80G GPUs.
+All experiments are conducted on $8\times$ 80G GPUs.
 For each query, we generate $2N=16$ rollouts.
@@
 On Geometry3K (Tab.
-2), our 3B model reaches 47.88% accuracy, outperforming GRPO by 5.2% and DAPO by 2.7%.; the 7B model achieves 55.89%, with gains of 3.3% and 1.4%, respectively.
+2), our 3B model reaches 47.88% accuracy, outperforming GRPO by 5.2% and DAPO by 2.7%; the 7B model achieves 55.89%, with gains of 3.3% and 1.4%, respectively.
 On out-of-domain math reasoning tasks, our method further improves average accuracy by 1.96% (3B) and 1.5% (7B) over GRPO, also surpassing DAPO.
@@
 Table 6: Extension experiments on Qwen2.5-VL-32B.
-Model MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
+Setting MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
 Qwen2.5-32B 57.0 38.2 75.4 72.9 71.3 80.7 65.9
@@
 Table 7: Extension experiments on Referring Expression Comprehension task.
-Model RefCOCO(testA) RefCOCO(testB) RefCOCO+(testA) RefCOCO+(testB) RefCOCOg(test)
+Setting RefCOCO(testA) RefCOCO(testB) RefCOCO+(testA) RefCOCO+(testB) RefCOCOg(test)
 Qwen2.5-3B 86.09 75.64 81.71 66.93 72.39
@@
 […]
-Scaling to 32B Model
+Scaling to 32B Model.
 We conduct experiments on Qwen2.5-VL-32B to inveritage the scalability and generalizability of Shuffle-R1 on larger models.
@@
 7 demonstrate that Shuffle-R1 also performs well on 32B scale.
-Referring Expression Comprehension
+Referring Expression Comprehension.
 For tasks beyond math/visual reasoning, we conduct an experiment on Referring Expression Comprehension (REC).
@@
 […]
-Language-only Reasoning
+Language-only Reasoning.
 While our initial analysis focused on the RL training dynamics of MLLMs, we conduct an early extension experiment to validate the feasibility of Shuffle-R1 on LLMs.
-We train Qwen2.5-Math-1.5B/7B-Base (Yang et al., 2024) on Open-S1 dataset (Dang and Ngo, 2025) for 150 steps and evaluate model performance on Math12K (Hendrycks et al., 2021), AIME24, MATH500 (Lightman et al., 2023), GSM8K (Cobbe et al., 2021), GPQA-Diamond (Rein et al., 2023) and OlympiadBench (He et al., 2024).
+We train Qwen2.5-Math-1.5B/7B-Base (Yang et al., 2024) on Open-S1 dataset (Dang and Ngo, 2025) for 150 steps and evaluate model performance on Math12K (Hendrycks et al., 2021), AIME24, MATH500 (Lightman et al., 2023), GSM8K (Cobbe et al., 2021), GPQA-Diamond (Rein et al., 2024) and OlympiadBench (He et al., 2024).
 As shown in Tab.
@@
 This work was done during the research internship of Linghao Zhu, Yiran Guan, and Dingkang Liang at Xiaomi.
-This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62225603 and 623B2038.
+This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62225603, 62576147 and 623B2038.
 Ethics statement
@@
 Computing Resources.
-All experiments are conducted on $8\times$ NVIDIA H800-80G-SXM GPUs.
+All experiments are conducted on $8\times$ 80G GPUs.
 Appendix E Dataset Collection
@@
 To fully evaluate model’s reasoning ability, apart from in-domain test set from training dataset, we select several representative benchmarks to examine model’s performance on math reasoning, visual perception, and chart understanding tasks.
-For math reasoning task, we choose MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024) and MathVista (Lu et al., 2023).
+For math reasoning task, we choose MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2025) and MathVista (Lu et al., 2023).
 We choose HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022) for visual perception and chart understanding tasks respectively.